### v1

# Sample-Based Conservative Bias Linear Power Flow Approximations

Abstract
The power flow equations are central to many problems in power system planning, analysis, and control. However, their inherent non-linearity and non-convexity present substantial challenges during problem-solving processes, especially for optimization problems. Accordingly, linear approximations are commonly employed to streamline computations, although this can often entail compromises in accuracy and feasibility. This paper proposes an approach termed Conservative Bias Linear Approximations (CBLA) for addressing these limitations. By minimizing approximation errors across a specified operating range while incorporating conservativeness (over- or under-estimating quantities of interest), CBLA strikes a balance between accuracy and tractability by maintaining linear constraints. By allowing users to design loss functions tailored to the specific approximated function, the bias approximation approach significantly enhances approximation accuracy. We illustrate the effectiveness of our proposed approach through several test cases.
Index Terms:
Conservative bias linear approximation; power flow approximation
I Introduction
The power flow equations play a central role in the operation and analysis of electrical power systems. These equations are essential for evaluating the behavior of power networks, making them key to various optimization problems such as resilient infrastructure planning [1, 2, 3], AC unit commitment [4, 5], and bilevel problems [6, 7]. However, the nonlinearity of the power flow equations induces non-convexities in these problems that pose significant computational challenges.
To address these challenges, researchers have developed various linear approximations such as DC power flow [8], LinDistFlow [9], first-order Taylor expansions of the power flow equations, and other approximations [10]. These methods offer simplified representations of power flow, which improve the tractability of power systems optimization problems. However, these linearizations often depend on broad assumptions such as maintaining voltages at 1 per unit and keeping voltage angle differences small between neighboring buses, as in DC power flow. These assumptions may not be valid across all operating conditions, potentially resulting in inaccuracies in the approximations. Consequently, the solutions derived from these linearized models may not closely align with the actual optimal solutions in real-world scenarios. This trade-off between simplicity and accuracy necessitates careful consideration when applying these linearizations in practice.
In response to these challenges, various studies have explored adaptive power flow approximations tailored to specific systems and operating ranges to enhance approximation accuracy (e.g., optimization-based approaches in [11] and sample-based approaches in [12, 13, 14]); see [15, 16, 17] for recent survey papers on this concept. For sample-based approaches, samples of operating points computed by repeatedly solving the power flow equations at various points within an operating range (e.g., a specified range of power injections) are leveraged to compute the approximations. By capturing complex nonlinear relationships directly from these samples, the resulting linear approximations can be more accurate as they are formulated by minimizing deviations from the solutions provided by the AC power flow equations for a particular system and operating range of interest. These adaptive power flow linearizations spend computational effort in computing the linearization coefficients in order to improve accuracy and tractability when applied in optimization problems. By trading up-front computational time for increased accuracy when applied, adaptive power flow approximations are particularly valuable in settings with both offline and online aspects where the linearization coefficients can be computed offline in advance of a real-time problem as well as settings where explicitly modeling power flow nonlinearities would lead to intractability, e.g., [1, 2, 3, 4, 5, 6, 7].
Extending the concept of sample-based adaptive power flow linearizations, the conservative linear approximation (CLA) approach in [12, 13] incorporates the concept of conservativeness. In other words, the CLAs are computed to minimize approximation errors with respect to the AC power flow equations while consistently over- or under-estimating quantities of interest over the set of drawn samples. The resulting approximations are particularly well suited for settings with an asymmetry in the implications of overestimating a quantity like voltage magnitude or current flow as opposed to underestimating that quantity. This is particularly relevant in power system optimization problems where feasibility is of paramount importance. For instance, when used in the bound on the magnitude of current flow through a line, a linearization that erroneously underestimates the amount of current flow risks predicting feasibility when the constraint is actually violated with respect to the nonlinear AC power flow equations. This is a more problematic linearization error than an overestimate of the current flow for use in this constraint. Thus, conservative linearizations that avoid errors in a particular direction (i.e., avoid either overestimates or underestimates of some quantity) are valuable in many power system optimization contexts. However, maintaining conservativeness can sometimes lead to reduced accuracy.
In this paper, we introduce an approach to approximating power flow equations called conservative bias linear approximation (CBLA). The CBLA approach seeks to balance the trade-off between conservativeness and accuracy, particularly in scenarios where certain samples are challenging to approximate accurately. To construct CBLAs, the process shares similarities with CLAs by beginning with drawing samples from within the operational range. These samples form the basis of a regression problem, which is solved to compute an approximated function representing the power flow equations. However, unlike CLA, CBLA does not explicitly enforce conservativeness in its approximated function as a hard constraint. Instead, the CBLA approach introduces an error function that penalizes linearization errors for samples that violate conservativeness to enable more accurate approximations.
CBLA offers the advantage of flexibility in designing customized error functions that quantify the penalty for deviating from actual values. User-defined error functions enable the approach to be tailored to particular quantities of interest and system characteristics, thus computing a linearization specialized for a specific problem. This flexibility can be particularly beneficial in scenarios where some violations are permissible, such as in chance-constrained optimization problems.
In summary, the main contributions of this paper are:
  1. (i)
    A CBLA formulation that is tailored to the specific system and operating range, optimal with respect to an error metric, and strikes a balance between conservativeness and accuracy.
  2. (ii)
    A discussion on choosing an error function for computing the CBLA.
  3. (iii)
    Numerical benchmarking of CBLAs for a variety of test cases.
The remainder of this paper is organized as follows: Section II covers background material on the power flow equations as well as sample-based conservative linear approximations. Section III introduces the conservative bias linear approximation approach. Section IV provides numerical results of our approach. Section V concludes the paper along with directions for our future work.
II Background
In this section, we provide background information about the AC power flow equations and present the recently developed conservative linear approximations of these equations.
II-A The Power Flow Equations
Consider a power system where a reference bus has the voltage angle set to 00. Let $V(\theta)$ denote the voltage magnitude (phasor). Let $P(Q)$ denote the active (reactive) power injection. We use the subscript $(\cdot)_{i}$ to represent a quantity at bus $i$ and the subscript $(\cdot)_{ik}$ to represent a quantity from or connecting bus $i$ to $k$. Let $j=\sqrt{-1}$. The AC power flow equations at bus $i$ are:
$P_{i}$ $=V_{i}^{2}G_{ii}+\sum_{k\in\mathcal{B}_{i}}V_{i}V_{k}(G_{ik}\cos\theta_{ik}+B_{ik}\sin\theta_{ik}),$ (1a)
$Q_{i}$ $=-V_{i}^{2}B_{ii}+\sum_{k\in\mathcal{B}_{i}}V_{i}V_{k}(G_{ik}\sin\theta_{ik}-B_{ik}\cos\theta_{ik}),$ (1b)
where $\theta_{ik}:=\theta_{i}-\theta_{k}$, $G(B)$ is a real (imaginary) part of the admittance matrix associated with the system, and $\mathcal{B}_{i}$ represents the set of all neighboring buses to bus $i$ including bus $i$ itself.
II-B Conservative linear approximations
The nonlinearity of the power flow equations in (1) contribute to the complexity encountered in solving optimization problems. To address this challenge, we previously introduced a sample-based conservative linear approximation (CLA) approach aimed at either over- or under-estimating specified quantities of interest, such as the magnitudes of voltages and current flows (as illustrated in Fig. 1)[12]. Moreover, CLAs facilitate parallel computation by enabling concurrent computation of the CLA for each quantity of interest. The construction of a CLA entails sampling power injections across an operational range of interest, such as a range of loads and power generated by Distributed Energy Resources (DERs), followed by computing power flow solutions for each sample and solving a constrained-regression problem.
For instance, samples for load demands are acquired utilizing a predefined probability distribution $\mathbb{P}_{\mathcal{S}}$ over a specified operational range $\mathcal{S}$. This range could be defined as $\mathcal{S}=\{P_{L_{d}}^{\text{min}}\leq P_{L_{d}}\leq P_{L_{d}}^{\text{max}},\enspace Q_{L_{d}}^{\text{min}}\leq Q_{L_{d}}\leq Q_{L_{d}}^{\text{max}}\mbox{for all }d\in\mathcal{N_{D}}\}$ where $(\,\cdot\,)_{L_{d}}$, where $(\,\cdot\,)_{L_{d}}$ denotes the load demand, $\mathbb{P}_{\mathcal{S}}$ represents the uniform distribution, and the superscripts max (min) indicate upper (lower) limits.
Figure 1: An illustration showcasing a comparison between a conventional linear approximation (on the left) and CLAs (on the right). In this visual representation, the solid line signifies the nonlinear function under consideration. In the figure on the left, the dotted line represents a traditional first-order Taylor approximation centered at point $\times$, while in the figure on the right, the dotted line above (below) corresponds to an over- (under-)estimating approximation.
The utilization of CLAs allows for the customization of the approximation to fit a specific operating range and the targeted system. Additionally, the sample-based approach enables integrating the behavior of complicated devices like tap-changing transformers and smart inverters into the approximation, as discussed in our prior work [6]. In the realm of optimization, CLAs offer a crucial advantage: they enable the satisfaction of nonlinear constraints while enforcing only linear inequalities, assuming the CLAs maintain conservativeness. Consequently, CLAs streamline optimization problems, rendering them suitable for commercial optimization solvers.
Consider a quantity of interest denoted as $\gamma$, which could represent variables such as the voltage magnitude at a specific bus or the magnitude of current flow along a particular line. In this context, bold quantities signify matrices and vectors. Let superscript $T$ denote the transpose. An overestimating CLA can be expressed as follows:
$a_{0}+\bm{a}_{1}^{T}\begin{bmatrix}\bm{P}\\ \bm{Q}\end{bmatrix}$ (2)
where $a_{0}$ is a scalar and $\bm{a}_{1}$ is a vector, both serving as decision variables in the regression problem later described in (4). This CLA is constructed to ensure the fulfillment of the following relationship for power injections $\bm{P}$ and $\bm{Q}$ within a specified range:
$\gamma\leq a_{0}+\bm{a}_{1}^{T}\begin{bmatrix}\bm{P}\\ \bm{Q}\end{bmatrix}.$ (3)
Assuming that (3) is indeed satisfied, we can ensure that the constraint $\gamma\leq\gamma^{\text{max}}$ is also satisfied by instead enforcing a linear constraint $a_{0}+\bm{a}_{1}^{T}\begin{bmatrix}\bm{P}\\ \bm{Q}\end{bmatrix}\leq\gamma^{\text{max}}$. This approach allows us to meet the upper bound requirement $\gamma^{\text{max}}$ without introducing the implicit system of nonlinear AC power flow equations in (1). Importantly, by employing the CLA, we are able to satisfy the nonlinear equations while maintaining a linear formulation, thus enhancing computational tractability without sacrificing feasibility in the resulting solution.
To compute a CLA, we solve for the coefficients of the affine function of power injections in (2) in the following regression problem:
$\min_{a_{0},\ \bm{a}_{1}}\quad\frac{1}{M}\sum_{m=1}^{M}\mathcal{L}\left(\gamma_{m}-\left(a_{0}+\bm{a}_{1}^{T}\begin{bmatrix}\bm{P}_{m}\\ \bm{Q}_{m}\end{bmatrix}\right)\right)$ (4a)
[…]
The subscript $(\,\cdot\,)_{m}$ denotes the $m^{\text{th}}$ sample and $M$ is the number of samples. The function $\mathcal{L}(\,\cdot\,)$ represents a loss function, such as the absolute value for $\ell_{1}$ loss or the square for squared-$\ell_{2}$ loss. In this paper, our focus is on quantities of interest denoted by $\gamma$, which correspond to the magnitudes of voltages ($V$) and current flows ($I$). The construction of underestimating CLAs follows a similar process as described in (4), with the key distinction being the reversal of the inequality direction in (4b).
The conservativeness of the CLA computed in (4) comes at the cost of reduced accuracy relative to the approximation corresponding to the unconstrained regression problem resulting from dropping (4a) from (4). To manage this tradeoff, the next section presents the main contribution of this paper, namely, a linear approximation technique that achieves a balance between conservativeness and accuracy. This approach involves biasing the linearization towards conservativeness, guided by a designated loss function.
III Conservative Bias Linear Approximations
The CLA approach presented in Section II-B is consistently conservative within the set of drawn samples. However, in certain scenarios, the conservativeness property may lead to significant errors due to specific samples. In this paper, we present a sample-based conservative bias linear approximation (CBLA) approach that is adaptive, meaning it can be tailored to a specific system and operating range. The CBLA approach is designed to be optimal, aiming to minimize a specific error metric while retaining a tendency to be conservative in order to enhance accuracy. This implies that the CBLA primarily minimizes errors between the approximating function and the samples, permitting samples to violate conservativeness at a specified cost.
III-A Formulation
Let $\epsilon$ denote the mismatch between the approximated quantity and the actual quantity. The optimization problem to compute a CBLA is formulated as follows:
$\min_{f(\epsilon_{m}(a_{0},\bm{a}_{1}))}\quad\frac{1}{M}\sum_{m=1}^{M}f(\epsilon_{m})$ (5)
where
$[\forall m=1,\ldots,M]$
$\epsilon_{m}=\gamma_{m}-\left(a_{0}+\bm{a}_{1}^{T}\begin{bmatrix}\bm{P}_{m}\\ \bm{Q}_{m}\end{bmatrix}\right),$ (6)
and
$f(\epsilon_{m})$ =$=\begin{dcases}g(\epsilon_{m}),&\text{if}\;\epsilon_{m}\leq 0\\ h(\epsilon_{m}),&\text{otherwise.}\\ \end{dcases}$ (7)
The optimization problem in (5) seeks to minimize the aggregated value of the error function $f(\,\cdot\,)$ defined in (7) over all samples by computing the coefficients $a_{0}$ and $\bm{a_{1}}$ in (6). This error function is contingent upon the error mismatch, denoted as $\epsilon$ in (6), between the estimated quantity and the actual quantity ($\gamma$). The error function is computed based on the sign of the error for each sample. In cases of overestimation, the value of $h(\epsilon_{m})$ is designed to be high, imposing a substantial cost for violation. Conversely, the value of $g(\epsilon_{m})$ is intended to be relatively low, reflecting scenarios where samples do not violate the conservativeness. Vice-versa, the value of $h(\epsilon_{m})$ is designed to be relatively low and $g(\epsilon_{m})$ to be high for an underestimation.
III-B Error function
Choosing a suitable error function in the CBLA approach is an important consideration. The choice of error function depends on various factors, such as the specific system requirements, a quantity of interest, and the trade-off between accuracy and conservativeness. To better understand how the error function works, we compare the error function used in the CLA approach with that of the CBLA approach.
The CLA approach imposes conservativeness across the set of sampled data in the constraints. We can rewrite the regression problem described in (4), which utilizes the $\ell_{1}$ loss function, as an optimization problem formulated in (5)–(7). In this formulation, the error function is defined as follows:
$f(\epsilon_{m})$ =$=\begin{dcases}\epsilon_{m},&\text{if}\;\epsilon_{m}\leq 0\\ \infty,&\text{otherwise.}\\ \end{dcases}$ (8)
Figure 2: An error function of the CLA where $g(\epsilon)=\epsilon$ and $h(\epsilon)=\infty$.
The error function defined in equation (8) (see Fig. 2) assigns an infinite cost to any violation of the overestimating requirement. This implies that for all samples drawn, no violation is permitted.
Figure 3: An example of an error function where $g(\epsilon)=\epsilon$ and $h(\epsilon)=e^{2\epsilon}-1$.
In contrast, our CBLA approach offers the flexibility to configure an error function that accommodates violations for specific samples, all while considering a predefined cost associated with these violations. This ability to tailor the error function empowers us to strike a balance between accuracy and the acceptable level of conservativeness for the specific system and operating range of interest. In Fig. 3, an example of an error function $f(\epsilon)$ for an overestimating CBLA is depicted. The function $f(\epsilon)$ exhibits a high value for $\epsilon>0$, $f(\epsilon)$ maintains a relatively low value for $\epsilon<0$, and $f(0)$ is zero, indicating an exact approximation.
In this setup, by assigning a higher penalty when an approximation violates conservativeness, we incline towards more conservative approximations while sacrificing some accuracy. Additionally, when the derivative of the function $h(\epsilon)$ increases as shown in Fig. 3, this function likely permits only small positive values of $\epsilon$, as larger positive values result in exponentially higher penalties. Conversely, reducing the cost tends to yield more accurate approximations at the expense of conservativeness.
When the error function is defined as a piecewise linear function (i.e., when both $g(\epsilon)$ and $h(\epsilon)$ are linear functions), the problem formulation to compute CBLAs in (5)–(7) can be framed as a linear program as follows:
$\min_{a_{0},\ \bm{a}_{1}}\quad\frac{1}{M}\sum_{m=1}^{M}z_{m}$ (9a)
[…]
where $z$ is a slack variable, and $k_{1}$ and $k_{2}$ are the coefficients of the linear error functions, i.e., $g(\epsilon)=k_{1}\epsilon$ and $h(\epsilon)=k_{2}\epsilon$.
With nonlinear error functions, the regression problem is a mixed-integer nonlinear program; nevertheless, it can be conveniently implemented using a user-defined function in Julia and subsequently solved with a Julia package like Optim, which provides a framework for solving constrained optimization problems [18].
IV Numerical results
In this section, we conduct numerical experiments on several test cases to examine the behavior of CBLA, highlight the benefits of error function design, and demonstrate the effectiveness of CBLA in a simplified OPF problem.
The test cases used in the simulations are case6ww, case14, and the IEEE 24-bus system, all of which are accessible in Matpower [19]. For approximations of voltage and current flow, we draw $500$ samples by varying the power injections within a range of $70\%$ to $130\%$ of their nominal values. Both voltage and current flow values are reported in per unit (pu). We use the $\ell_{1}$ norm as the loss function $\mathcal{L}(\,\cdot\,)$. The numerical simulation was conducted in Julia using the Optim package.
IV-A Conservative Bias Linear Approximations
We begin our numerical tests by examining the effects of changing the error function in (7) in our CBLA approach. As discussed in Section III-B, error functions are designed to balance conservativeness and accuracy. By testing different error functions, we aim to understand how they impact the number of violated samples and accuracy of the approximated power flow equations.
(a) $g(\epsilon)=\epsilon^{2},h(\epsilon)=\epsilon^{2}$.
(b) $g(\epsilon)=\epsilon^{2},h(\epsilon)=10\epsilon^{2}$.
(c) $g(\epsilon)=\epsilon^{2},h(\epsilon)=100\epsilon^{2}$.
(d) $g(\epsilon)=\epsilon^{2},h(\epsilon)=1000\epsilon^{2}$.
Figure 4: Plots of results when (a) $\alpha=1$ (equivalent to the squared-$\ell_{2}$ loss), (b) $\alpha=10$, (c) $\alpha=100$, and (d) $\alpha=1000$ for current flow from bus 3 to bus 24 in IEEE 24-bus system. The red points represent overestimating CBLAs. The black line represents a zero approximation error.
In Fig. 4, we present the results of using the CBLA approach to intentionally overestimate the predicted current flow from bus 3 to bus 24 in IEEE 24-bus system using different quadratic error functions. The error functions employed in this test are defined as $g(\epsilon)=x^{2}$ and $h(\epsilon)=\alpha x^{2}$, where $\alpha$ is a parameter that we vary across the test. Specifically, we adjust $\alpha$ to take values of 1, 10, 100, and 1000. When $\alpha=1$, the error functions are equivalent (i.e., $g(\epsilon)=h(\epsilon)$), indicating that there is no difference in cost between violating and not violating conservativeness (i.e., the error function is the squared-$\ell_{2}$ loss). Under this condition, almost all samples are well approximated, but several samples fall below the zero approximation error line, indicating a deviation from the overestimating objective.
As we increase $\alpha$, fewer samples fall below the zero approximation error line, suggesting improved adherence to the overestimation goal. However, this improvement comes at the cost of lower overall accuracy. At the highest value of $\alpha=1000$, most samples are overestimated as intended, but the level of accuracy appears to be the lowest among all the plots. This trade-off highlights the importance of carefully selecting the value of $\alpha$ to achieve a suitable balance between overestimation and accuracy.
Figure 5: Results showing the average error per sample in per unit (pu) and the number of violated samples due to overestimating CBLA of current flow from bus 3 to bus 24 in IEEE 24-bus system, as the value of $\alpha$ (labeled at each point) varies from $1$ to $10^{4}$.
To gain further insight into the effects of varying $\alpha$, we plot the relationship between the average error per sample of the approximated flow and the number of violated samples when varying the value of $\alpha$ in Fig. 5. In this test, we adjust $\alpha$ over a range from $1$ to $10^{4}$. The results reveal a clear trend: as $\alpha$ increases, the average error per sample also increases while the number of violated samples decreases significantly, demonstrating the trade-off between conservativeness and accuracy. This is due to the increased enforcement of conservativeness in the error function. Specifically, the average error per sample increases from $0.00869$ when $\alpha=1$ to $0.0455$ when $\alpha=10^{4}$, while the number of violated samples decreases from $161$ when $\alpha=1$ to just $12$ when $\alpha=10^{4}$.
TABLE I: Approximated current flow errors and number of violated samples at specific buses in IEEE 24-bus system
Line Average errors/sample # violated samples
(From-to) $\alpha=1$ $\alpha=10^{2}$ $\alpha=10^{4}$ $\alpha=1$ $\alpha=10^{2}$ $\alpha=10^{4}$
3-14 0.00869 0.03012 0.04551 161 25 12
6-10 0.00907 0.02274 0.03780 202 30 8
9-12 0.01621 0.04961 0.09397 180 29 6
The data presented in Table I illustrates the relationship between the number of violated samples and the average approximated current flow errors across different values of $\alpha$ at different lines. These results align with the trend observed in Fig. 5, confirming that as $\alpha$ increases, the average error per sample increases while the number of violated samples decreases.
IV-B Application: Simplified optimal power flow
While our main goal is to use our CBLA approach for challenge problems such as bilevel problems and mixed-integer nonlinear programs, these are outside the scope of this paper. Instead, we concentrate on showcasing results within a simplified optimal power flow (OPF) framework, which offers a conceptual demonstration and a foundation for comparing different linear approximations (DC power flow and conservative bias linear approximation (CBLA)). This simplified version of the OPF scenario imposes constraints on voltages at buses where $P$ and $Q$ are known, and power generation within defined ranges, excluding line flow limits (see [12] for the full problem setup).
TABLE II: Results from OPF comparing solutions from AC-, DC-, and CBLA-OPF
Formulation Case
case6ww case14
AC-OPF 2986.04 5368.30
DC-OPF 2995.15 (0.31%) 5368.52 (0.004%)
Violation $V$ (0.029 pu) No violation
CBLA-OPF ($\alpha=1$) 2987.28 (0.04%) 5368.52 (0.004%)
Violation No violation $V$ (0.004 pu)
CBLA-OPF ($\alpha=10^{2}$) 2987.42 (0.05%) 5368.52 (0.004%)
Violation No violation No violation
CBLA-OPF ($\alpha=10^{6}$) 2987.51 (0.05%) 5368.52 (0.004%)
Violation No violation No violation
Table II presents the outcomes of applying the AC power flow equations and various power flow approximations in a simplified optimal power flow (OPF) scenario. The AC-OPF serves as the baseline, with no violations and providing a reference for comparing the percentage difference in optimal cost against other formulations. In the overestimating CBLA approach, we utilize error functions $g(\epsilon)=\epsilon^{2}$ and $h(\epsilon)=\alpha\epsilon^{2}$, while for underestimating CBLA, we use error functions $g(\epsilon)=\alpha\epsilon^{2}$ and $h(\epsilon)=\epsilon^{2}$. As discussed in Section IV-A, setting $\alpha=1$ indicates that the error function does not differentiate between overestimating or underestimating the linearization error.
The results indicate that the DC-OPF leads to a maximum voltage violation of $0.029$ pu and an optimal cost that is $0.31\%$ higher than the AC-OPF’s solution for case6ww. In case14, the CBLA-OPF causes only a minor voltage violation of $0.004$ pu when $\alpha=1$. When $\alpha=10^{2}$ and $10^{6}$ (with the latter closely representing the conservative linear approximations in our previous work [12]), both test cases have no voltage violations. Moreover, the optimal costs using DC-OPF and CBLA-OPF are only $0.004\%$ higher than the optimal solution from AC-OPF in case14. These results suggest that the flexibility of CBLA allows us to select a suitable function that optimizes cost depending on the specific system being analyzed.
V Conclusion and future work
This paper presents a conservative bias linear approximation (CBLA) approach for approximating the power flow equations. This approach strives to balance conservativeness and accuracy while maintaining linearity in the approximations. The numerical results highlight the potential advantages of using CBLA for power flow problems. By selecting an appropriate error function, we can achieve an effective balance between conservativeness and accuracy. Additionally, the ability to choose different error functions allows CBLA to be tailored to specific systems and operational conditions, ultimately enhancing performance and reliability in power system optimization.
In our future work, we aim to extend our current approach by developing additional conservative bias approximations through the use of piecewise linearizations formulated as neural networks. Moreover, we plan to apply our proposed approach to a broader range of power system planning and resilience tasks. This includes tackling complex bilevel problems as well as conducting capacity expansion planning studies.
Acknowledgement
The authors thank the Department of Mechatronics Engineering at Rajamangala University of Technology Phra Nakhon.

### v2

@@
-Sample-Based Conservative Bias Linear Power Flow Approximations
+Paprapee Buason, Sidhant Misra Los Alamos National Laboratory Los Alamos, NM, USA
 Abstract
@@
 By allowing users to design loss functions tailored to the specific approximated function, the bias approximation approach significantly enhances approximation accuracy.
-We illustrate the effectiveness of our proposed approach through several test cases.
+We illustrate the effectiveness of our proposed approach through several test cases, including its application to a unit commitment problem, where CBLA consistently achieves lower operating costs and improved feasibility compared to traditional linearization methods.
 Index Terms:
-Conservative bias linear approximation; power flow approximation
+Conservative bias linear approximation; power flow approximation; unit commitment.
 I Introduction
-The power flow equations play a central role in the operation and analysis of electrical power systems.
+The power flow equations play a central role in the operation and analysis of electric power systems.
 These equations are essential for evaluating the behavior of power networks, making them key to various optimization problems such as resilient infrastructure planning [1, 2, 3], AC unit commitment [4, 5], and bilevel problems [6, 7].
@@
 These methods offer simplified representations of power flow, which improve the tractability of power systems optimization problems.
-However, these linearizations often depend on broad assumptions such as maintaining voltages at 1 per unit and keeping voltage angle differences small between neighboring buses, as in DC power flow.
+However, these linearizations often depend on broad assumptions such as maintaining voltages near 1 per unit and keeping voltage angle differences small between neighboring buses, as in DC power flow.
 These assumptions may not be valid across all operating conditions, potentially resulting in inaccuracies in the approximations.
@@
 By capturing complex nonlinear relationships directly from these samples, the resulting linear approximations can be more accurate as they are formulated by minimizing deviations from the solutions provided by the AC power flow equations for a particular system and operating range of interest.
-These adaptive power flow linearizations spend computational effort in computing the linearization coefficients in order to improve accuracy and tractability when applied in optimization problems.
+These adaptive power flow linearizations spend computational effort to calculate the linearization coefficients in order to improve accuracy and tractability when applied in optimization problems.
 By trading up-front computational time for increased accuracy when applied, adaptive power flow approximations are particularly valuable in settings with both offline and online aspects where the linearization coefficients can be computed offline in advance of a real-time problem as well as settings where explicitly modeling power flow nonlinearities would lead to intractability, e.g., [1, 2, 3, 4, 5, 6, 7].
@@
 This flexibility can be particularly beneficial in scenarios where some violations are permissible, such as in chance-constrained optimization problems.
-In summary, the main contributions of this paper are:
+This work extends our previous study presented in [18], where we introduced the CBLA approach for power flow linearization.
+In this journal version, we build upon that foundation by incorporating a unit commitment formulation [19] using CBLA, comparing it against other power flow linearization methods, and detailing a step-by-step process for solving the unit commitment problem.
+In summary, the paper’s main contributions are:
   1. (i)
@@
   3. (iii)
-    Numerical benchmarking of CBLAs for a variety of test cases.
-The remainder of this paper is organized as follows: Section II covers background material on the power flow equations as well as sample-based conservative linear approximations.
-Section III introduces the conservative bias linear approximation approach.
-Section IV provides numerical results of our approach.
-Section V concludes the paper along with directions for our future work.
+    Numerical tests of CBLAs for a variety of test cases.
+  4. (iv)
+    An application and numerical results on proposed unit commitment problems comparing the CBLA approach with other power flow linearizations.
+The remainder of this paper is organized as follows: Section II presents background material on power flow equations and sample-based conservative linear approximations.
+Section III introduces the proposed conservative bias linear approximation (CBLA) approach.
+Section IV describes the iterative method for solving the unit commitment problem using CBLA.
+Section V provides numerical results demonstrating the performance of the proposed method.
+Finally, Section VI concludes the paper and outlines directions for future research.
 II Background
@@
 Consider a power system where a reference bus has the voltage angle set to 00.
-Let $V(\theta)$ denote the voltage magnitude (phasor).
+Let $V(\theta)$ denote the voltage magnitude (angle).
 Let $P(Q)$ denote the active (reactive) power injection.
@@
 $Q_{i}$ $=-V_{i}^{2}B_{ii}+\sum_{k\in\mathcal{B}_{i}}V_{i}V_{k}(G_{ik}\sin\theta_{ik}-B_{ik}\cos\theta_{ik}),$ (1b)
-where $\theta_{ik}:=\theta_{i}-\theta_{k}$, $G(B)$ is a real (imaginary) part of the admittance matrix associated with the system, and $\mathcal{B}_{i}$ represents the set of all neighboring buses to bus $i$ including bus $i$ itself.
+where $\theta_{ik}:=\theta_{i}-\theta_{k}$, $G$ and $B$ are the real and imaginary parts of the system’s admittance matrix, respectively, and $\mathcal{B}_{i}$ is the set of all neighboring buses to bus $i$, including bus $i$.
 II-B Conservative linear approximations
-The nonlinearity of the power flow equations in (1) contribute to the complexity encountered in solving optimization problems.
+The nonlinearity of the power flow equations in (1) contributes to the complexity encountered in solving optimization problems.
 To address this challenge, we previously introduced a sample-based conservative linear approximation (CLA) approach aimed at either over- or under-estimating specified quantities of interest, such as the magnitudes of voltages and current flows (as illustrated in Fig.
@@
 For instance, samples for load demands are acquired utilizing a predefined probability distribution $\mathbb{P}_{\mathcal{S}}$ over a specified operational range $\mathcal{S}$.
-This range could be defined as $\mathcal{S}=\{P_{L_{d}}^{\text{min}}\leq P_{L_{d}}\leq P_{L_{d}}^{\text{max}},\enspace Q_{L_{d}}^{\text{min}}\leq Q_{L_{d}}\leq Q_{L_{d}}^{\text{max}}\mbox{for all }d\in\mathcal{N_{D}}\}$ where $(\,\cdot\,)_{L_{d}}$, where $(\,\cdot\,)_{L_{d}}$ denotes the load demand, $\mathbb{P}_{\mathcal{S}}$ represents the uniform distribution, and the superscripts max (min) indicate upper (lower) limits.
+This range could be defined as $\mathcal{S}=\{P_{L_{d}}^{\text{min}}\leq P_{L_{d}}\leq P_{L_{d}}^{\text{max}},\enspace Q_{L_{d}}^{\text{min}}\leq Q_{L_{d}}\leq Q_{L_{d}}^{\text{max}}\mbox{for all }L_{d}\in\mathcal{N_{D}}\}$, where $(\,\cdot\,)_{L_{d}}$ denotes the load demand, $\mathbb{P}_{\mathcal{S}}$ represents the uniform distribution, $\mathcal{N_{D}}$ is the set of all buses with loads, and the superscripts max (min) indicate upper (lower) limits.
 Figure 1: An illustration showcasing a comparison between a conventional linear approximation (on the left) and CLAs (on the right).
@@
 where $a_{0}$ is a scalar and $\bm{a}_{1}$ is a vector, both serving as decision variables in the regression problem later described in (4).
-This CLA is constructed to ensure the fulfillment of the following relationship for power injections $\bm{P}$ and $\bm{Q}$ within a specified range:
+This CLA is designed to ensure the following relationship for power injections $\bm{P}$ and $\bm{Q}$ within a specified range:
 $\gamma\leq a_{0}+\bm{a}_{1}^{T}\begin{bmatrix}\bm{P}\\ \bm{Q}\end{bmatrix}.$ (3)
@@
 In this paper, we present a sample-based conservative bias linear approximation (CBLA) approach that is adaptive, meaning it can be tailored to a specific system and operating range.
-The CBLA approach is designed to be optimal, aiming to minimize a specific error metric while retaining a tendency to be conservative in order to enhance accuracy.
-This implies that the CBLA primarily minimizes errors between the approximating function and the samples, permitting samples to violate conservativeness at a specified cost.
+The CBLA approach is designed to be optimal, minimizing a specific error metric while maintaining a tendency to be conservative to enhance accuracy.
+This means CBLA focuses on minimizing errors between the approximating function and the samples, allowing some samples to violate conservativeness at a specified cost.
 III-A Formulation
@@
 This error function is contingent upon the error mismatch, denoted as $\epsilon$ in (6), between the estimated quantity and the actual quantity ($\gamma$).
-The error function is computed based on the sign of the error for each sample.
-In cases of overestimation, the value of $h(\epsilon_{m})$ is designed to be high, imposing a substantial cost for violation.
-Conversely, the value of $g(\epsilon_{m})$ is intended to be relatively low, reflecting scenarios where samples do not violate the conservativeness.
-Vice-versa, the value of $h(\epsilon_{m})$ is designed to be relatively low and $g(\epsilon_{m})$ to be high for an underestimation.
+The error function is computed based on the error’s sign for each sample.
+For overestimation, $h(\epsilon_{m})$ is set high to impose a substantial cost for violation, while $g(\epsilon_{m})$ is kept low for non-violations.
+Conversely, for underestimation, $h(\epsilon_{m})$ is relatively low and $g(\epsilon_{m})$ is high.
+For an alternative formulation of CBLA that incorporates a penalty term to influence the optimization behavior, see the appendix.
 III-B Error function
@@
 Figure 2: An error function of the CLA where $g(\epsilon)=\epsilon$ and $h(\epsilon)=\infty$.
-The error function defined in equation (8) (see Fig. 2) assigns an infinite cost to any violation of the overestimating requirement. This implies that for all samples drawn, no violation is permitted.
+The error function in (8) (see Fig. 2) assigns an infinite cost to any violation of the overestimating requirement. This implies that for all samples drawn, no violation is permitted.
 Figure 3: An example of an error function where $g(\epsilon)=\epsilon$ and $h(\epsilon)=e^{2\epsilon}-1$.
@@
 The function $f(\epsilon)$ exhibits a high value for $\epsilon>0$, $f(\epsilon)$ maintains a relatively low value for $\epsilon<0$, and $f(0)$ is zero, indicating an exact approximation.
-In this setup, by assigning a higher penalty when an approximation violates conservativeness, we incline towards more conservative approximations while sacrificing some accuracy.
-Additionally, when the derivative of the function $h(\epsilon)$ increases as shown in Fig.
-3, this function likely permits only small positive values of $\epsilon$, as larger positive values result in exponentially higher penalties.
-Conversely, reducing the cost tends to yield more accurate approximations at the expense of conservativeness.
+In this setup, assigning a higher penalty for violations of conservativeness leads to more conservative approximations at the expense of some accuracy.
+Additionally, when the derivative of the function $h(\epsilon)$ increases, as shown in Fig.
+3, the function tends to allow only small positive values of $\epsilon$, because larger positive values result in exponentially higher penalties.
+Conversely, reducing the penalty cost tends to produce more accurate approximations but with less conservativeness.
 When the error function is defined as a piecewise linear function (i.e., when both $g(\epsilon)$ and $h(\epsilon)$ are linear functions), the problem formulation to compute CBLAs in (5)–(7) can be framed as a linear program as follows:
@@
 where $z$ is a slack variable, and $k_{1}$ and $k_{2}$ are the coefficients of the linear error functions, i.e., $g(\epsilon)=k_{1}\epsilon$ and $h(\epsilon)=k_{2}\epsilon$.
-With nonlinear error functions, the regression problem is a mixed-integer nonlinear program; nevertheless, it can be conveniently implemented using a user-defined function in Julia and subsequently solved with a Julia package like Optim, which provides a framework for solving constrained optimization problems [18].
-IV Numerical results
-In this section, we conduct numerical experiments on several test cases to examine the behavior of CBLA, highlight the benefits of error function design, and demonstrate the effectiveness of CBLA in a simplified OPF problem.
-The test cases used in the simulations are case6ww, case14, and the IEEE 24-bus system, all of which are accessible in Matpower [19].
+Nonlinear error functions transform the regression problem into a mixed-integer nonlinear program.
+This can be efficiently implemented using a user-defined function in Julia and solved with packages such as Optim, which offers a robust framework for constrained optimization problems [20].
+IV Unit Commitment
+The unit commitment (UC) problem is an optimization problem in power system operations that aims to determine the optimal schedule of generating units over a specified time horizon, typically 24 hours.
+The primary goal is to minimize the total cost of operation while ensuring that the demand for electricity is met and the technical and operational constraints of the generation units and transmission network are satisfied.
+The variability and uncertainty associated with loads, particularly when load shedding may be necessary, and generation introduce additional complexities that the unit commitment (UC) problem must address to ensure grid reliability and operational efficiency.
+As a mixed-integer optimization problem, the UC formulation involves binary variables to represent the on/off status of generators, along with continuous variables to model power output.
+IV-A Computational procedure and parallelization potential
+Figure 4: A flowchart illustrating the computational process for solving the unit commitment (UC) problem using CBLA.
+Steps enclosed in red boxes correspond to CBLA, while steps enclosed in blue boxes relate to the UC problem.
+Steps marked with * indicate parallelizable processes.
+Solving the unit commitment problem using CBLA involves multiple computational steps, each contributing to accurately determining generator commitments and dispatch while leveraging efficient approximations of power flow equations.
+Given the complexity of unit commitment and the nonlinearity of AC power flow, CBLA serves as an effective method to construct computationally efficient linear approximations.
+The overall process consists of three key steps:
+  1. 1.
+    CBLA computation: Solve the CBLA formulation in (5) to approximate voltage and current over a range of operating conditions.
+This step requires generating samples within expected operating ranges and solving the CBLA optimization problem.
+  2. 2.
+    Unit commitment optimization: Using the CBLA approximations obtained in Step 1, solve the UC problem to determine generator statuses for each generator at each hour while ensuring operational constraints are satisfied.
+  3. 3.
+    AC optimal power flow (AC OPF) evaluation: Given the generator statuses from Step 2, solve the AC OPF that allows load shedding to determine the generation dispatch at each hour, ensuring feasibility under AC power flow constraints [21].
+This computational framework allows parallel execution, particularly in Steps 1 and 3.
+In Step 1, solving CBLA involves two key computations that can be parallelized.
+First, the power flow solutions at each operating point can be computed independently.
+Second, the regression problem for the CBLA can be solved independently for each quantity, using the same set of samples across all regression problems.
+Similarly, in Step 3, since ramping constraints are not enforced and only minimum up- and down-time constraints are considered, the AC OPF computations at different time steps are independent once generator statuses are fixed, enabling parallel execution across time periods.
+Fig.
+4 illustrates the full set of computational steps required to solve the unit commitment problem using the CBLA approach.
+IV-B Optimization Problem Formulation
+Let $u_{g,t}$ be a generator on/off status and $s_{g,t}\;(r_{g,t})$ be a generator startup (shutdown) status for generator $g$ at time $t$.
+The unit commitment problem with the DC power flow [22] is formulated as follows:
+$\min$ $\sum_{t\in\mathcal{T}}\left(\sum_{g\in\mathcal{G}}c_{g}(P_{g,t})+C_{g}^{\text{startup}}s_{g,t}+C_{g}^{\text{shutdown}}r_{g,t}\right)$ (10a)
+s.t. $\sum_{g\in\mathcal{G}}P_{g,t}=D_{t},\quad\forall t\in\mathcal{T},$ (10b)
+$[\forall g\in\mathcal{G}]$
+$P_{g,t}\leq P_{g}^{\text{max}}u_{g,t},\;P_{g,t}\geq P_{g}^{\text{min}}u_{g,t},\;\forall t\in\mathcal{T},$ (10c)
+$s_{g,t}-r_{g,t}=u_{g,t}-u_{g,t-1},\;\forall t\in\mathcal{T},$ (10d)
+$\sum_{\tau=t}^{t+T_{g}^{\text{up}}-1}u_{g,\tau}\geq T_{g}^{\text{up}}\big{(}u_{g,t}-u_{g,t-1}\big{)},$
+$\forall t\in\{1,\ldots,T-T_{g}^{\text{up}}\},$ (10e)
+$\sum_{\tau=t}^{t+T_{g}^{\text{down}}-1}(1-u_{g,\tau})\geq T_{g}^{\text{down}}\big{(}u_{g,t-1}-u_{g,t}\big{)},$
+$\forall t\in\{1,\ldots,T-T_{g}^{\text{down}}\},$ (10f)
+$\theta_{i,t}-\theta_{k,t}=\frac{P_{ik,t}}{B_{ik}},\;\forall(i,k)\in\mathcal{L},\forall t\in\mathcal{T},$ (10g)
+$-P_{ik}^{\text{max}}\leq P_{ik,t}\leq P_{ik}^{\text{max}},\;\forall(i,k)\in\mathcal{L},\forall t\in\mathcal{T},$ (10h)
+where $u_{g,0}=0$ for all $g$, $c_{g}(P_{g,t})$ is a cost associated with generator $g$ at time $t$ (for example, $c_{g}(P_{g,t})=c_{2g}P_{g,t}^{2}+c_{1g}P_{g,t}+c_{0g}$ for a polynomial cost function), $\mathcal{G}$ is a set of all generators, $\mathcal{L}$ is a set of all lines, and $\mathcal{T}$ is a set of time in hours.
+The optimization problem in (10) is referred to as the UC problem, which minimizes the total operating cost over a given time horizon.
+The objective function in (10a) comprises generation costs, startup costs, and shutdown costs.
+The constraint (10b) ensures power balance at every time step, requiring total generation to match load demand, where (10c) enforces that the power output of each generator stays within its operational limits when the generator is online.
+The constraint (10d) models the relationship between the on/off status of generators and their startup/shutdown states.
+In this setup, we consider ramping constraints as part of the startup and shutdown constraints, ensuring that generation changes smoothly during transitions.
+The constraints (10e)–(10f) enforce minimum up-time and down-time requirements for generators, ensuring they remain online or offline for a specified duration after being switched on or off.
+The constraint (10g) represents the linearized DC power flow equations, relating power flows to voltage angle differences and line susceptance.
+Finally, (10h) ensures that power flow on each line does not exceed its thermal limits in either direction.
+To extend this formulation and provide a baseline for comparison, the appendix presents a set of linear constraints based on the first-order Taylor approximation of voltage magnitudes and current flows.
+IV-C Unit commitment using conservative bias linear approximations
+The unit commitment problem formulated in the previous subsection enforces network feasibility using a DC power flow approximation.
+While computationally efficient, the DC power flow model does not explicitly incorporate voltage and current constraints.
+In this section, we introduce penalty terms and replace the DC power flow constraints in (10g)–(10h) with a CBLA, which provides a linearized yet conservative representation of these constraints.
+Let $\mathcal{B}$ be a set of all buses. The CBLA enforces voltage magnitude limits at each bus using the following inequalities:
+$\underline{a}_{0,i,t}+\underline{\bm{a}}_{1,i,t}^{T}\begin{bmatrix}\bm{P_{t}}\\ \bm{Q_{t}}\end{bmatrix}\geq V_{i}^{\text{min}}-v_{i,t}^{\text{min}},\;\forall i\in\mathcal{B},\forall t\in\mathcal{T},$ (11a)
+$\overline{a}_{0,i,t}+\overline{\bm{a}}_{1,i,t}^{T}\begin{bmatrix}\bm{P_{t}}\\ \bm{Q_{t}}\end{bmatrix}\leq V_{i}^{\text{max}}+v_{i,t}^{\text{max}},\;\forall i\in\mathcal{B},\forall t\in\mathcal{T},$ (11b)
+where $v_{i,t}^{\text{min}}$ and $v_{i,t}^{\text{max}}$ are nonnegative penalty variables that allow violation of the voltage approximations at bus $i$ at time $t$.
+Here, $\bm{P}_{t}$ ($\bm{Q}_{t}$) denotes the vector of active (reactive) power injections at time $t$, while $V_{i}^{\text{min}}$ and $V_{i}^{\text{max}}$ are the lower and upper voltage magnitude limits at bus $i$.
+The coefficients $\underline{a}_{0,i}$ and $\underline{\bm{a}}_{1,i}$ correspond to the underestimating voltage approximation at bus $i$, and $\overline{a}_{0,i,t}$ and $\overline{\bm{a}}_{1,i,t}$ correspond to the overestimating voltage approximation.
+Additionally, we enforce current flow limits in transmission lines through the following constraint:
+$\overline{b}_{0,ik,t}+\overline{\bm{b}}_{1,ik,t}^{T}\begin{bmatrix}\bm{P_{t}}\\ \bm{Q_{t}}\end{bmatrix}\leq I_{ik}^{\text{max}}+z_{ik,t},\;\forall(i,k)\in\mathcal{L},\forall t\in\mathcal{T},$ (12)
+where $z_{ik,t}$ is a nonnegative penalty variable allowing violation of the current flow approximation on line $(i,k)$ at time $t$, $I_{ik,t}^{\text{max}}$ denotes the maximum allowable current on line $(i,k)$, and $\overline{b}_{0,ik,t}$, $\overline{\bm{b}}_{1,ik,t}$ are coefficients for the overestimating current flow.
+Let $F(\bm{P}_{g,t},\bm{s}_{g,t},\bm{r}_{g,t})$ denote the objective function in (10a). With the inclusion of penalty terms, the modified objective becomes:
+$F(\bm{P},\bm{s},\bm{r})+\sum_{t\in\mathcal{T}}\left(\lambda\sum_{(i,k)\in\mathcal{L}}z_{ik,t}+\gamma\sum_{i\in\mathcal{B}}(v_{i,t}^{\text{max}}+v_{i,t}^{\text{min}})\right),$ (13)
+where $\lambda$ and $\gamma$ are the penalty weights associated with violations of current and voltage magnitude constraints, respectively.
+The CBLA method constructs linear constraints from coefficients derived at sampled operating points within a specified range.
+Before solving the UC problem, we can identify buses and lines that do not exhibit voltage or current violations in the sampled data; constraints for these elements remain inactive and can be excluded, reducing computational complexity without compromising feasibility.
+By incorporating these constraints into the UC problem, CBLA ensures generation schedules remain feasible with respect to voltage and line flow limits while maintaining computational tractability.
+Unlike the standard DC power flow model, CBLA systematically accounts for these limits through conservative yet efficient linear approximations.
+Its sample-based nature also enables constraint screening, identifying likely binding constraints and allowing flexibility in selecting the error function depending on whether voltage or line flow limits are active.
+To maintain the validity of CBLA-based constraints during the UC process, we adopt an iterative approach (Fig.
+4).
+In each iteration, we solve the UC problem with the current CBLA, then check if the generation statuses and dispatch levels match those from a first-order Taylor-based method or previous CBLA iterations.
+We also verify whether power injections stay within the sampled region used to build the CBLA.
+If violations occur, we identify the affected time periods and buses or lines, draw new samples around the updated solution, recompute the CBLA, and resolve the UC problem.
+This process prevents accuracy degradation from extrapolation and cycling between inconsistent UC solutions.
+This CBLA-based formulation further allows two practical enhancements.
+First, it offers flexibility in controlling conservativeness by adjusting a bias term when constructing the constraints, enabling the user to trade off between feasibility margin and conservatism.
+Second, it supports the inclusion of penalty terms that allow for controlled constraint violations, enhancing feasibility in cases where a strictly conservative formulation may be overly restrictive.
+IV-D Combinatorial Growth of Valid Generator Statuses
+In the unit commitment formulation, each generator’s on/off status over time must satisfy minimum up and down time constraints (10e)–(10f), substantially reducing the number of admissible on/off sequences.
+Nevertheless, the feasible status space still grows exponentially with the number of time periods, making brute-force enumeration impractical.
+To quantify this space, we implement an enumeration algorithm that counts binary vectors of length $\lvert\mathcal{T}\rvert$ satisfying a minimum run length $d=T^{\text{up}}=T^{\text{down}}$ for both states, allowing exceptions for an initial short off-run and final short runs.
+Due to these exceptions, a clean closed-form expression is difficult to derive.
+Instead, the number of valid vectors is characterized recursively: sequences are built by appending valid runs (of at least length $d$) to shorter valid sequences, naturally leading to a dynamic programming approach.
+Let $f(\lvert\mathcal{T}\rvert,s,\ell)$ denote the number of valid binary sequences of length $\lvert\mathcal{T}\rvert$, ending in state $s\in\{0,1\}$ with a current run of length $\ell$.
+The recurrence relation is:
+$f(\lvert\mathcal{T}\rvert,s,\ell)=\begin{cases}\sum\limits_{k=d}^{\lfloor\nicefrac{{\lvert\mathcal{T}\rvert}}{{2}}\rfloor}f(\lvert\mathcal{T}\rvert-k,1-s,k),\;\text{if }\ell=d,\lvert\mathcal{T}\rvert\geq d\\[4.30554pt] f(\lvert\mathcal{T}\rvert-1,s,\ell-1),\;\text{if }\ell>d\\[4.30554pt] 1,\;\text{if }\lvert\mathcal{T}\rvert=\ell\leq d\text{, one of the exceptions holds}\\[4.30554pt] 0,\;\text{otherwise}\end{cases}$
+where the exceptions include:
+  • Initial 0’s from position 1 with length less than $d$,
+  • Final 0’s or 1’s from position $\lvert\mathcal{T}\rvert-d+2$ onward.
+The total number of valid sequences is
+$F(\lvert\mathcal{T}\rvert)=\sum_{s\in\{0,1\}}\sum_{\ell=1}^{\lvert\mathcal{T}\rvert}f(\lvert\mathcal{T}\rvert,s,\ell).$
+For example, with $\lvert\mathcal{T}\rvert=5$ and $d=2$, applying the recurrence yields 13 valid sequences for a single generator.
+Therefore, for 6 generators, the total number of configurations is $(F(5))^{6}=13^{6}=4,\!826,\!809$, demonstrating how the feasible status space grows rapidly even for modest systems.
+This growth illustrates why brute-force enumeration of all valid generator commitment combinations quickly becomes intractable.
+To manage this complexity, we avoid exhaustive search by solving the unit commitment problem iteratively: each iteration only considers a manageable subset of decisions, and additional constraints are introduced in subsequent iterations to guide the solution away from unstable commitment patterns, such as cycling or oscillating generator decisions, i.e., repeatedly looping through statuses.
+The algorithm is explained in more detail in Section IV-A.
+IV-E AC Optimal Power Flow with Load Shedding
+The AC Optimal Power Flow (AC OPF) problem determines the optimal generation dispatch by solving the full nonlinear AC power flow equations while satisfying network constraints.
+Using generator commitment statuses from the unit commitment problem in (10) (solved with formulations such as DC power flow, first-order Taylor approximations, or CBLAs), we compute the dispatch for each online generator at each hour.
+The multi-period formulation minimizes the total cost across all hours while ensuring power balance and network feasibility at each time.
+If constraints prevent meeting full demand, load shedding is used as a last resort, with $L_{i,t}$ denoting the load shed at bus $i$ and time $t$.
+The AC OPF with load shedding is formulated as follows:
+$\min$ $\sum_{t\in\mathcal{T}}\left(\sum_{g\in\mathcal{G}}c_{g}(P_{g,t})+\sum_{i\in\mathcal{B}}C^{\text{shed}}(L_{i,t})\right)$ (14a)
+s.t. $[\forall i\in\mathcal{B},\forall t\in\mathcal{T},\forall g\in\mathcal{G},\forall(i,k)\in\mathcal{L}]$
+$P_{i,t}^{\text{gen}}-P_{i,t}^{\text{load}}+L_{i,t}=$
+$\quad\sum_{k\in\mathcal{B}}V_{i,t}V_{k,t}\big{(}G_{ik}\cos\theta_{ik,t}+B_{ik}\sin\theta_{ik,t}\big{)},$ (14b)
+$Q_{i,t}^{\text{gen}}-Q_{i,t}^{\text{load}}+\beta_{i}L_{i,t}=$
+$\quad\sum_{k\in\mathcal{B}}V_{i,t}V_{k,t}\big{(}G_{ik}\sin\theta_{ik,t}-B_{ik}\cos\theta_{ik,t}\big{)},$ (14c)
+$P_{g,t}^{\text{min}}u_{g,t}\leq P_{g,t}\leq P_{g,t}^{\text{max}}u_{g,t},$ (14d)
+$Q_{g,t}^{\text{min}}u_{g,t}\leq Q_{g,t}\leq Q_{g,t}^{\text{max}}u_{g,t},$ (14e)
+$V_{i,t}^{\text{min}}\leq V_{i,t}\leq V_{i,t}^{\text{max}},$ (14f)
+$S_{ik,t}=\sqrt{P_{ik,t}^{2}+Q_{ik,t}^{2}}\leq S_{ik}^{\text{max}},$ (14g)
+$0\leq L_{i,t}\leq P_{i,t}^{\text{load}},\quad 0\leq\beta_{i}L_{i,t}\leq Q_{i,t}^{\text{load}},$ (14h)
+where $\beta_{i}=\tan\alpha_{i}$, given that $\cos\alpha_{i}$ represents the power factor associated with load $i$.
+The objective function in (14a) minimizes the total cost of generation over all time periods while penalizing load shedding using a function $C^{\text{shed}}(L_{i,t})$ to discourage demand curtailment unless necessary.
+The power balance constraints in (14b)–(14c) enforce the nonlinear AC power flow equations for active and reactive power at each bus and time period.
+The generator limits in (14d)–(14e) ensure that active and reactive power generation respects the unit commitment status $u_{g,t}$.
+The voltage limits in (14f) ensure that bus voltages remain within safe operating ranges.
+The thermal constraints in (14g) restrict apparent power flows on transmission lines to their rated capacities.
+Finally, the load shedding constraints in (14h) ensure that the curtailed demand remains non-negative and does not exceed the original load at any bus and time period.
+This formulation integrates the unit commitment results into AC OPF, optimizing the dispatch of committed generators while allowing for load shedding with constant power factors.
+V Numerical results
+In this section, we conduct numerical experiments on several test cases to examine the behavior of CBLA, highlight the benefits of error function design, and demonstrate the effectiveness of CBLA in a UC problem.
+The test cases used in the simulations are the IEEE 24-bus system and the IEEE 30-bus system, all of which are accessible in Matpower [23].
 For approximations of voltage and current flow, we draw $500$ samples by varying the power injections within a range of $70\%$ to $130\%$ of their nominal values.
@@
 We use the $\ell_{1}$ norm as the loss function $\mathcal{L}(\,\cdot\,)$.
-The numerical simulation was conducted in Julia using the Optim package.
-IV-A Conservative Bias Linear Approximations
+The numerical simulation was conducted in MATLAB using Yalmip [24] and in Julia using the Optim package.
+V-A Conservative Bias Linear Approximations
 We begin our numerical tests by examining the effects of changing the error function in (7) in our CBLA approach.
 As discussed in Section III-B, error functions are designed to balance conservativeness and accuracy.
-By testing different error functions, we aim to understand how they impact the number of violated samples and accuracy of the approximated power flow equations.
+By testing various error functions, we aim to understand their impact on the number of violated samples and the accuracy of the approximated power flow equations.
 (a) $g(\epsilon)=\epsilon^{2},h(\epsilon)=\epsilon^{2}$.
@@
 In Fig.
-4, we present the results of using the CBLA approach to intentionally overestimate the predicted current flow from bus 3 to bus 24 in IEEE 24-bus system using different quadratic error functions.
-The error functions employed in this test are defined as $g(\epsilon)=x^{2}$ and $h(\epsilon)=\alpha x^{2}$, where $\alpha$ is a parameter that we vary across the test.
-Specifically, we adjust $\alpha$ to take values of 1, 10, 100, and 1000.
-When $\alpha=1$, the error functions are equivalent (i.e., $g(\epsilon)=h(\epsilon)$), indicating that there is no difference in cost between violating and not violating conservativeness (i.e., the error function is the squared-$\ell_{2}$ loss).
-Under this condition, almost all samples are well approximated, but several samples fall below the zero approximation error line, indicating a deviation from the overestimating objective.
-As we increase $\alpha$, fewer samples fall below the zero approximation error line, suggesting improved adherence to the overestimation goal.
-However, this improvement comes at the cost of lower overall accuracy.
-At the highest value of $\alpha=1000$, most samples are overestimated as intended, but the level of accuracy appears to be the lowest among all the plots.
+5, we present the results of using the CBLA approach to intentionally overestimate the predicted current flow from bus 3 to bus 24 in IEEE 24-bus system using different quadratic error functions.
+The error functions used in this test are defined as $g(\epsilon)=\epsilon^{2}$ and $h(\epsilon)=\alpha\epsilon^{2}$, where $\alpha$ is a parameter that we vary across the test.
+Specifically, the parameter $\alpha$ modifies the bias in the CBLA formulation; a higher value of $\alpha$ increases the tendency toward conservativeness.
+We adjust $\alpha$ to take values of 1, 10, 100, and 1000.
+When $\alpha=1$, the error functions are equivalent (i.e., $g(\epsilon)=h(\epsilon)$), implying no distinction in cost between violating and not violating conservativeness (i.e., the squared-$\ell_{2}$ loss).
+In this scenario, most samples are well approximated, but some fall below the zero approximation error line, indicating a deviation from the overestimating objective.
+Increasing $\alpha$ reduces the number of samples falling below the zero approximation error line, indicating better adherence to the overestimation goal.
+However, this improvement sacrifices overall accuracy.
+At $\alpha=1000$, where most samples are intentionally overestimated, the accuracy tends to be lower compared to other scenarios.
 This trade-off highlights the importance of carefully selecting the value of $\alpha$ to achieve a suitable balance between overestimation and accuracy.
@@
 To gain further insight into the effects of varying $\alpha$, we plot the relationship between the average error per sample of the approximated flow and the number of violated samples when varying the value of $\alpha$ in Fig.
-5.
+6.
 In this test, we adjust $\alpha$ over a range from $1$ to $10^{4}$.
@@
 Specifically, the average error per sample increases from $0.00869$ when $\alpha=1$ to $0.0455$ when $\alpha=10^{4}$, while the number of violated samples decreases from $161$ when $\alpha=1$ to just $12$ when $\alpha=10^{4}$.
-TABLE I: Approximated current flow errors and number of violated samples at specific buses in IEEE 24-bus system
+TABLE I: Approximated current flow errors and number of violated samples at representative buses in IEEE 24-bus system
 Line Average errors/sample # violated samples
@@
 9-12 0.01621 0.04961 0.09397 180 29 6
-The data presented in Table I illustrates the relationship between the number of violated samples and the average approximated current flow errors across different values of $\alpha$ at different lines.
+The data in Table I illustrates the relationship between the number of violated samples and the average approximated current flow errors across different $\alpha$ values at different lines.
 These results align with the trend observed in Fig.
-5, confirming that as $\alpha$ increases, the average error per sample increases while the number of violated samples decreases.
-IV-B Application: Simplified optimal power flow
-While our main goal is to use our CBLA approach for challenge problems such as bilevel problems and mixed-integer nonlinear programs, these are outside the scope of this paper.
-Instead, we concentrate on showcasing results within a simplified optimal power flow (OPF) framework, which offers a conceptual demonstration and a foundation for comparing different linear approximations (DC power flow and conservative bias linear approximation (CBLA)).
-This simplified version of the OPF scenario imposes constraints on voltages at buses where $P$ and $Q$ are known, and power generation within defined ranges, excluding line flow limits (see [12] for the full problem setup).
-TABLE II: Results from OPF comparing solutions from AC-, DC-, and CBLA-OPF
-Formulation Case
-case6ww case14
-AC-OPF 2986.04 5368.30
-DC-OPF 2995.15 (0.31%) 5368.52 (0.004%)
-Violation $V$ (0.029 pu) No violation
-CBLA-OPF ($\alpha=1$) 2987.28 (0.04%) 5368.52 (0.004%)
-Violation No violation $V$ (0.004 pu)
-CBLA-OPF ($\alpha=10^{2}$) 2987.42 (0.05%) 5368.52 (0.004%)
-Violation No violation No violation
-CBLA-OPF ($\alpha=10^{6}$) 2987.51 (0.05%) 5368.52 (0.004%)
-Violation No violation No violation
-Table II presents the outcomes of applying the AC power flow equations and various power flow approximations in a simplified optimal power flow (OPF) scenario.
-The AC-OPF serves as the baseline, with no violations and providing a reference for comparing the percentage difference in optimal cost against other formulations.
-In the overestimating CBLA approach, we utilize error functions $g(\epsilon)=\epsilon^{2}$ and $h(\epsilon)=\alpha\epsilon^{2}$, while for underestimating CBLA, we use error functions $g(\epsilon)=\alpha\epsilon^{2}$ and $h(\epsilon)=\epsilon^{2}$.
-As discussed in Section IV-A, setting $\alpha=1$ indicates that the error function does not differentiate between overestimating or underestimating the linearization error.
-The results indicate that the DC-OPF leads to a maximum voltage violation of $0.029$ pu and an optimal cost that is $0.31\%$ higher than the AC-OPF’s solution for case6ww.
-In case14, the CBLA-OPF causes only a minor voltage violation of $0.004$ pu when $\alpha=1$.
-When $\alpha=10^{2}$ and $10^{6}$ (with the latter closely representing the conservative linear approximations in our previous work [12]), both test cases have no voltage violations.
-Moreover, the optimal costs using DC-OPF and CBLA-OPF are only $0.004\%$ higher than the optimal solution from AC-OPF in case14.
-These results suggest that the flexibility of CBLA allows us to select a suitable function that optimizes cost depending on the specific system being analyzed.
-V Conclusion and future work
-This paper presents a conservative bias linear approximation (CBLA) approach for approximating the power flow equations.
-This approach strives to balance conservativeness and accuracy while maintaining linearity in the approximations.
+6, confirming that as $\alpha$ increases, the average error per sample increases while the number of violated samples decreases.
+V-B Application: Unit commitment
+The comparison is conducted on the IEEE 30-bus test system, which has 6 generators.
+Voltage magnitude limits are set to 0.94 p.u.
+and 1.05 p.u.
+at all buses, and current magnitude limits are derived from the values in mpc.branch(:,RATE_A) in the MATPOWER case files.
+The cost of load shedding is $C^{\text{shed}}(L_{i,t})=500L_{i,t}$.
+We draw $1000$ samples by varying power injections, both loads and generations, to cover possible load shedding and generation variation within a range of $70\%$ to $130\%$ of their nominal values.
+As shown in Table II, we report both the total cost, which includes load shedding penalties, and the generation-only cost, which excludes them to better isolate the impact of load shedding.
+Although the DC-based formulation often yields lower generation-only costs, it leads to much higher total costs due to frequent and severe load shedding.
+As illustrated in Fig.
+7, the generator commitment schedules from CBLA differ from those of the DC and first-order Taylor methods, highlighting CBLA’s ability to adapt commitment decisions based on a more accurate power flow approximation.
+Figure 7: Generator commitment schedules for the 30-bus system using three different power flow linearization methods: DC, first-order Taylor, and CBLA (when $\alpha=10$ and $100$).
+Each row block corresponds to a generator across five scheduling hours.
+Blue indicates the generator is ON, while red indicates it is OFF.
+This difference is especially evident in hour 4, where the DC-based approach leads to 11.82% load shedding and a total cost of \$20,417.
+In comparison, the first-order Taylor approximation reduces load shedding to 5.01%, with a much lower cost of \$11,683.
+CBLA reduces load shedding even further to just 1.92%, while lowering the total cost to \$8,053.7—clearly demonstrating the value of its adaptive linearization.
+TABLE II: Hourly Comparison Across Different Approximations
+H1 H2 H3 H4 H5
+Load Demand (MW) 152.2 192.13 234.29 287.33 219.27
+DC Load Shedding (%) 0% 4.37% 2.57% 11.82% 15.23%
+Outputs (MW)
+Gen 1 158.28 193.44 159.05 205.66 196.71
+Gen 2 0 0 80 61.62 0
+Gen 3 – Gen 6 0 0 0 0 0
+Total Cost (\$) 1818 6328 6090 20417 18869
+Gen-Only Cost (\$) 1817.6 2135.2 3076.1 3431.5 2167.3
+First-Order Taylor Load Shedding (%) 0% 0% 0.70% 5.01% 2.42%
+Outputs (MW)
+Gen 1 52.84 85.27 134.71 202.56 189.72
+Gen 2 68.48 80 80 62.73 0
+Gen 3 0 0 0 0 0
+Gen 4 33.86 32.41 26.96 22.16 33.90
+Gen 5 – Gen 6 0 0 0 0 0
+Total Cost (\$) 3483 3682 4800 11683 5875
+Gen-Only Cost (\$) 3483.0 3682.1 3978.0 4480.5 3219.1
+CBLA ($\alpha=10$) Load Shedding (%) 0% 0% 0% 1.92% 0%
+Outputs (MW)
+Gen 1 58.14 58.32 99.31 152.81 127.40
+Gen 2 0 0 0 0 0
+Gen 3 26.96 27.47 45.73 46.77 50.00
+Gen 4 50.40 55.00 53.42 50.02 47.48
+Gen 5 18.73 24.61 0 0 0
+Gen 6 0 29.04 40 40 0
+Total Cost (\$) 4506.2 5660.4 4929.7 8053.7 3958.8
+Gen-Only Cost (\$) 4506.2 5660.4 4929.7 5299.5 3958.8
+CBLA ($\alpha=100$) Load Shedding (%) 0% 0% 0% 1.92% 0.93%
+Outputs (MW)
+Gen 1 110.53 110.45 99.31 152.81 100.66
+Gen 2 0 0 0 0 80.00
+Gen 3 45.01 45.40 45.73 46.77 42.51
+Gen 4 0 0 53.42 50.02 0
+Gen 5 0 0 0 0 0
+Gen 6 0 40.00 40.00 40.00 0
+Total Cost (\$) 2637.0 3799.1 4929.7 8053.7 4830.8
+Gen-Only Cost (\$) 2637.0 3799.1 4929.7 5299.5 3811.4
+While the DC and first-order Taylor methods rely on a single linearization of nonlinear power flows, CBLA builds a sample-based conservative approximation over a defined operating region, delivering significant improvements.
+Across all five hours, CBLA eliminates load shedding in four hours and limits it to minimal levels in the most challenging hour.
+CBLA consistently achieves the lowest total cost.
+For example, in hour 5, CBLA attains a lower total cost of \$3,958.8 compared to \$5,875 with the Taylor approximation and \$18,869 with the DC approximation.
+The table also includes results for a more conservative CBLA variant with $\alpha=100$.
+With $\alpha=100$, the UC solution tends to avoid constraint violations.
+However, this added conservativeness comes at the cost of reduced accuracy in approximations.
+As a result, when the generation decisions from the UC problem are passed to the AC OPF, the resulting solution may still exhibit load shedding—particularly in Hour 5, where 0.93% of the demand remains unmet.
+This trade-off underscores the need to balance conservativeness and approximation accuracy.
+CBLA’s tunable parameter $\alpha$ allows us to adjust this balance to achieve the desired feasibility and flexibility.
+By tailoring the approximation to system conditions, CBLA helps reduce constraint violations and yields UC solutions that generally remain feasible under the nonlinear AC power flow model.
+Moreover, the benefits of CBLA become more pronounced in larger systems or under tight feasibility margins.
+Unlike the first-order Taylor approximation, which is inherently local, CBLA is constructed to remain conservative across a sampled region of interest.
+This range-aware property improves cost-efficiency in decision-making.
+Compared to traditional methods, CBLA provides a more adaptive solution framework that maintains the tractability of linear programming while reducing the gap between the models of linear and nonlinear power systems.
+VI Conclusion and future work
+This paper presents a conservative bias linear approximation (CBLA) approach for approximating the power flow equations, aiming to balance conservativeness and accuracy while preserving linearity.
 The numerical results highlight the potential advantages of using CBLA for power flow problems.
-By selecting an appropriate error function, we can achieve an effective balance between conservativeness and accuracy.
-Additionally, the ability to choose different error functions allows CBLA to be tailored to specific systems and operational conditions, ultimately enhancing performance and reliability in power system optimization.
-In our future work, we aim to extend our current approach by developing additional conservative bias approximations through the use of piecewise linearizations formulated as neural networks.
+Selecting suitable error functions enables an effective balance between conservativeness and accuracy.
+Additionally, the ability to choose different error functions allows CBLA to be tailored to specific systems and operational conditions.
+When applied to the unit commitment problem, the CBLA-based formulation consistently outperformed formulations based on DC power flow and first-order Taylor approximations, achieving lower overall operating costs in the subsequent evaluation of the AC optimal power flow.
+In our future work, we aim to extend our current approach by developing additional conservative bias approximations through the use of piecewise linearizations.
 Moreover, we plan to apply our proposed approach to a broader range of power system planning and resilience tasks.
-This includes tackling complex bilevel problems as well as conducting capacity expansion planning studies.
+This includes tackling complex bilevel problems, conducting capacity expansion planning studies, and extending the application of CBLA to unbalanced three-phase systems.
 Acknowledgement
-The authors thank the Department of Mechatronics Engineering at Rajamangala University of Technology Phra Nakhon.
+P. Buason thanks the Department of Mechatronics Engineering at Rajamangala University of Technology Phra Nakhon.
+Appendix: First-Order Taylor Approximation of Voltage and Current Constraints
+Suppose the power injections $\bm{P}$ and $\bm{Q}$ deviate from their nominal values $\bm{P}^{\text{nom}}$ and $\bm{Q}^{\text{nom}}$.
+Let $\mathcal{B}^{\text{PQ}}$ denote the set of PQ buses, $\widetilde{V}_{i,t}$ denote the first-order Taylor approximation of the voltage magnitude at bus $i$ and time $t$, and let the abbreviation “nom” indicate that a quantity is evaluated at the nominal operating point, which is computed separately for each hour based on the corresponding loads.
+For each $i\in\mathcal{B}^{\text{PQ}}$, the first-order Taylor approximation of the voltage magnitude is:
+$\widetilde{V}_{i,t}=V_{i}^{\text{nom}}+\frac{\partial V_{i}}{\partial\bm{P}}\bigg{|}_{\text{nom}}\cdot(\bm{P}_{t}-\bm{P}^{\text{nom}})+\frac{\partial V_{i}}{\partial\bm{Q}}\bigg{|}_{\text{nom}}\cdot(\bm{Q}_{t}-\bm{Q}^{\text{nom}}),$ (15)
+where the partial derivatives $\frac{\partial V_{i}}{\partial\bm{P}}$ and $\frac{\partial V_{i}}{\partial\bm{Q}}$ are evaluated at the nominal operating point using the inverse of the power flow Jacobian.
+For non-PQ buses, the voltage magnitude is fixed at its nominal value:
+$\widetilde{V}_{i,t}=V_{i}^{\text{nom}},\quad\forall i\notin\mathcal{B}^{\text{PQ}}.$ (16)
+To model current flow limits, we approximate the real part of the power flow on each transmission line $(i,k)\in\mathcal{L}$ using the difference in bus voltage magnitudes from (16):
+$\Re(S_{ik,t})\approx|Y_{ik}|\cdot(\widetilde{V}_{i,t}-\widetilde{V}_{k,t}),$ (17)
+where $\Re(\;\cdot\;)$ denotes the real part of a complex quantity and $|Y_{ik}|$ is the magnitude of the admittance between buses $i$ and $k$.
+We then enforce the following absolute value inequality to approximate current magnitude limits:
+$-I_{ik}^{\text{max}}\leq|Y_{ik}|\cdot(\widetilde{V}_{i,t}-\widetilde{V}_{k,t})\leq I_{ik}^{\text{max}},\;\forall(i,k)\in\mathcal{L},\;\forall t\in\mathcal{T}.$ (18)
+To allow controlled violations of these approximations similar to (11)–(12), slack variables may be added as needed in the optimization formulation.

### v3

@@
-Paprapee Buason, Sidhant Misra Los Alamos National Laboratory Los Alamos, NM, USA
+Paprapee Buason12, Sidhant Misra1, Daniel K. Molzahn3
 Abstract
-The power flow equations are central to many problems in power system planning, analysis, and control.
-However, their inherent non-linearity and non-convexity present substantial challenges during problem-solving processes, especially for optimization problems.
-Accordingly, linear approximations are commonly employed to streamline computations, although this can often entail compromises in accuracy and feasibility.
-This paper proposes an approach termed Conservative Bias Linear Approximations (CBLA) for addressing these limitations.
-By minimizing approximation errors across a specified operating range while incorporating conservativeness (over- or under-estimating quantities of interest), CBLA strikes a balance between accuracy and tractability by maintaining linear constraints.
-By allowing users to design loss functions tailored to the specific approximated function, the bias approximation approach significantly enhances approximation accuracy.
-We illustrate the effectiveness of our proposed approach through several test cases, including its application to a unit commitment problem, where CBLA consistently achieves lower operating costs and improved feasibility compared to traditional linearization methods.
-Index Terms:
-Conservative bias linear approximation; power flow approximation; unit commitment.
+Accurate modeling of power flow behavior is essential for a wide range of power system applications, yet the nonlinear and nonconvex structure of the underlying equations often limits their direct use in large-scale optimization problems.
+As a result, linear models are frequently adopted to improve computational tractability, though these simplifications can introduce excessive approximation error or lead to constraint violations.
+This paper presents a linear approximation framework, referred to as Conservative Bias Linear Approximations (CBLA), that systematically incorporates conservativeness into the approximation process.
+Rather than solely minimizing local linearization error, CBLA constructs linear constraints that bound the nonlinear functions of interest over a defined operating region while reducing overall approximation bias.
+The proposed approach maintains the simplicity of linear formulations and allows the approximation to be shaped through user-defined loss functions tailored to specific system quantities.
+Numerical studies demonstrate that CBLA provides more reliable and accurate approximations than conventional linearization techniques, and its integration into a unit commitment formulation results in improved feasibility and reduced operating costs.
 I Introduction
-The power flow equations play a central role in the operation and analysis of electric power systems.
-These equations are essential for evaluating the behavior of power networks, making them key to various optimization problems such as resilient infrastructure planning [1, 2, 3], AC unit commitment [4, 5], and bilevel problems [6, 7].
-However, the nonlinearity of the power flow equations induces non-convexities in these problems that pose significant computational challenges.
-To address these challenges, researchers have developed various linear approximations such as DC power flow [8], LinDistFlow [9], first-order Taylor expansions of the power flow equations, and other approximations [10].
-These methods offer simplified representations of power flow, which improve the tractability of power systems optimization problems.
-However, these linearizations often depend on broad assumptions such as maintaining voltages near 1 per unit and keeping voltage angle differences small between neighboring buses, as in DC power flow.
-These assumptions may not be valid across all operating conditions, potentially resulting in inaccuracies in the approximations.
-Consequently, the solutions derived from these linearized models may not closely align with the actual optimal solutions in real-world scenarios.
-This trade-off between simplicity and accuracy necessitates careful consideration when applying these linearizations in practice.
-In response to these challenges, various studies have explored adaptive power flow approximations tailored to specific systems and operating ranges to enhance approximation accuracy (e.g., optimization-based approaches in [11] and sample-based approaches in [12, 13, 14]); see [15, 16, 17] for recent survey papers on this concept.
-For sample-based approaches, samples of operating points computed by repeatedly solving the power flow equations at various points within an operating range (e.g., a specified range of power injections) are leveraged to compute the approximations.
-By capturing complex nonlinear relationships directly from these samples, the resulting linear approximations can be more accurate as they are formulated by minimizing deviations from the solutions provided by the AC power flow equations for a particular system and operating range of interest.
-These adaptive power flow linearizations spend computational effort to calculate the linearization coefficients in order to improve accuracy and tractability when applied in optimization problems.
-By trading up-front computational time for increased accuracy when applied, adaptive power flow approximations are particularly valuable in settings with both offline and online aspects where the linearization coefficients can be computed offline in advance of a real-time problem as well as settings where explicitly modeling power flow nonlinearities would lead to intractability, e.g., [1, 2, 3, 4, 5, 6, 7].
-Extending the concept of sample-based adaptive power flow linearizations, the conservative linear approximation (CLA) approach in [12, 13] incorporates the concept of conservativeness.
-In other words, the CLAs are computed to minimize approximation errors with respect to the AC power flow equations while consistently over- or under-estimating quantities of interest over the set of drawn samples.
-The resulting approximations are particularly well suited for settings with an asymmetry in the implications of overestimating a quantity like voltage magnitude or current flow as opposed to underestimating that quantity.
-This is particularly relevant in power system optimization problems where feasibility is of paramount importance.
-For instance, when used in the bound on the magnitude of current flow through a line, a linearization that erroneously underestimates the amount of current flow risks predicting feasibility when the constraint is actually violated with respect to the nonlinear AC power flow equations.
-This is a more problematic linearization error than an overestimate of the current flow for use in this constraint.
-Thus, conservative linearizations that avoid errors in a particular direction (i.e., avoid either overestimates or underestimates of some quantity) are valuable in many power system optimization contexts.
-However, maintaining conservativeness can sometimes lead to reduced accuracy.
-In this paper, we introduce an approach to approximating power flow equations called conservative bias linear approximation (CBLA).
-The CBLA approach seeks to balance the trade-off between conservativeness and accuracy, particularly in scenarios where certain samples are challenging to approximate accurately.
-To construct CBLAs, the process shares similarities with CLAs by beginning with drawing samples from within the operational range.
-These samples form the basis of a regression problem, which is solved to compute an approximated function representing the power flow equations.
-However, unlike CLA, CBLA does not explicitly enforce conservativeness in its approximated function as a hard constraint.
-Instead, the CBLA approach introduces an error function that penalizes linearization errors for samples that violate conservativeness to enable more accurate approximations.
-CBLA offers the advantage of flexibility in designing customized error functions that quantify the penalty for deviating from actual values.
-User-defined error functions enable the approach to be tailored to particular quantities of interest and system characteristics, thus computing a linearization specialized for a specific problem.
-This flexibility can be particularly beneficial in scenarios where some violations are permissible, such as in chance-constrained optimization problems.
-This work extends our previous study presented in [18], where we introduced the CBLA approach for power flow linearization.
-In this journal version, we build upon that foundation by incorporating a unit commitment formulation [19] using CBLA, comparing it against other power flow linearization methods, and detailing a step-by-step process for solving the unit commitment problem.
-In summary, the paper’s main contributions are:
+Accurate representation of network physics is fundamental to a broad range of power system optimization problems.
+The nonlinear alternating current (AC) power flow equations govern voltage magnitudes, phase angles, and line flows, and therefore directly influence planning, operational scheduling, and reliability assessments.
+These equations appear in applications including resilient infrastructure design [gupta_buason_molzahn-fairness_pv_limit, owen_aquino_talkington_molzahn-EVacuation_feeder, bhusal2020], AC unit commitment (UC) [castillo2016, Sampath2018], and bilevel or market-based formulations [buason2023datadriven, wogrin2020].
+However, embedding the full nonlinear equations into optimization models introduces nonconvexities that significantly complicate computation, particularly in large-scale or mixed-integer settings.
+To mitigate these computational challenges, numerous linear approximations of the power flow equations have been developed.
+Widely used models such as DC power flow [stott2009], LinDistFlow [baran1989], and first-order Taylor expansions [wood2013power] provide tractable representations that enable efficient optimization.
+While these approximations are attractive due to their simplicity, they are typically derived under restrictive assumptions, for example, small voltage angle differences or near-nominal voltage magnitudes.
+When system operating conditions deviate from these assumptions, the resulting linear models may produce substantial errors, potentially leading to suboptimal or infeasible solutions when evaluated against the underlying AC equations.
+Thus, a fundamental trade-off arises between computational tractability and physical accuracy.
+Recognizing these limitations, recent research has investigated adaptive linearization strategies that tailor approximations to a specific system and operating region.
+Rather than relying on globally valid but coarse assumptions, these methods construct linear models using information obtained from representative operating samples.
+Optimization-based techniques [misra2018optimal] and sample-driven approaches [BUASON2022, buason2024adaptive, CHEN2022108573] compute linear coefficients by minimizing deviations from AC power flow solutions within a prescribed range of power injections; see [jia2023tutorial1, 10202779, jia2023tutorial2] for surveys of this emerging direction.
+By shifting computational effort to an offline stage where samples are generated and processed, adaptive approximations can provide improved fidelity while preserving linear structure for subsequent optimization tasks.
+Among these approaches, the conservative linear approximation (CLA) framework [BUASON2022, buason2024adaptive] incorporates an additional structural property: directional conservativeness.
+In CLA, the linear approximation is constructed so that, across the sampled operating region, quantities of interest are consistently overestimated or underestimated relative to the AC solution.
+This feature is particularly valuable in optimization problems where certain types of approximation error are more harmful than others.
+For example, underestimating line currents in a thermal limit constraint may falsely indicate feasibility, whereas overestimation results primarily in conservativeness.
+By enforcing one-sided error behavior, CLA reduces the risk of infeasible solutions.
+However, strict conservativeness may increase overall approximation error, especially in regions where the nonlinear relationship is difficult to represent accurately with a single linear model.
+Motivated by this observation, this paper introduces a generalized framework termed conservative bias linear approximation (CBLA).
+The CBLA methodology retains the sample-based construction process but relaxes strict one-sided enforcement of conservativeness.
+Instead of imposing conservativeness as a hard constraint, CBLA incorporates it through a bias-aware error function that penalizes violations asymmetrically.
+This formulation enables a controlled balance between directional bias and approximation accuracy.
+As a result, CBLA can reduce excessive conservativeness while still discouraging the most critical forms of approximation error.
+A distinguishing feature of CBLA is its flexibility in error function design.
+Users may define loss functions that reflect the operational priorities of a given application, allowing the approximation to emphasize particular quantities or error directions.
+This flexibility makes the framework adaptable to diverse optimization contexts, including those in which limited constraint violations are tolerable, such as chance-constrained or risk-aware formulations.
+By tailoring the regression objective to the needs of the downstream problem, CBLA produces linear models that are specialized to both the system and the intended optimization task.
+This work extends our preliminary study in [buason_cbla], where the CBLA concept was first introduced for power flow approximation.
+In the present journal version, we expand the framework and demonstrate its practical implications through extensive numerical evaluation.
+In particular, we integrate CBLA into a UC formulation [Padhy2004], compare its performance with alternative linearization techniques, and describe a structured solution procedure for solving the resulting problem.
+The results illustrate that CBLA achieves improved feasibility and reduced operating cost relative to conventional linear models.
+The primary contributions of this paper are summarized as follows:
   1. (i)
-    A CBLA formulation that is tailored to the specific system and operating range, optimal with respect to an error metric, and strikes a balance between conservativeness and accuracy.
+    A CBLA framework that balances directional conservativeness and approximation accuracy over a specified operating range.
   2. (ii)
-    A discussion on choosing an error function for computing the CBLA.
+    A discussion of error function selection and its role in shaping the trade-off between bias and fidelity.
   3. (iii)
-    Numerical tests of CBLAs for a variety of test cases.
+    Numerical validation of CBLA across multiple test systems.
   4. (iv)
-    An application and numerical results on proposed unit commitment problems comparing the CBLA approach with other power flow linearizations.
-The remainder of this paper is organized as follows: Section II presents background material on power flow equations and sample-based conservative linear approximations.
-Section III introduces the proposed conservative bias linear approximation (CBLA) approach.
-Section IV describes the iterative method for solving the unit commitment problem using CBLA.
-Section V provides numerical results demonstrating the performance of the proposed method.
-Finally, Section VI concludes the paper and outlines directions for future research.
-II Background
-In this section, we provide background information about the AC power flow equations and present the recently developed conservative linear approximations of these equations.
-II-A The Power Flow Equations
-Consider a power system where a reference bus has the voltage angle set to 00.
-Let $V(\theta)$ denote the voltage magnitude (angle).
-Let $P(Q)$ denote the active (reactive) power injection.
-We use the subscript $(\cdot)_{i}$ to represent a quantity at bus $i$ and the subscript $(\cdot)_{ik}$ to represent a quantity from or connecting bus $i$ to $k$.
-Let $j=\sqrt{-1}$.
-The AC power flow equations at bus $i$ are:
-$P_{i}$ $=V_{i}^{2}G_{ii}+\sum_{k\in\mathcal{B}_{i}}V_{i}V_{k}(G_{ik}\cos\theta_{ik}+B_{ik}\sin\theta_{ik}),$ (1a)
-$Q_{i}$ $=-V_{i}^{2}B_{ii}+\sum_{k\in\mathcal{B}_{i}}V_{i}V_{k}(G_{ik}\sin\theta_{ik}-B_{ik}\cos\theta_{ik}),$ (1b)
-where $\theta_{ik}:=\theta_{i}-\theta_{k}$, $G$ and $B$ are the real and imaginary parts of the system’s admittance matrix, respectively, and $\mathcal{B}_{i}$ is the set of all neighboring buses to bus $i$, including bus $i$.
+    An application of CBLA to UC, including comparative results against established linear power flow models.
+The remainder of this paper is organized as follows.
+Section II reviews relevant background on power flow modeling and conservative linear approximations.
+Section III presents the proposed CBLA methodology.
+Section IV details the UC formulation and solution approach.
+Section V provides numerical results, and Section VI concludes the paper.
+II Background on AC Power Flow and Conservative Linear Approximations
+This section reviews the AC power flow model and summarizes the sample-based conservative linear approximation (CLA) framework that motivates the development of the proposed approach.
+II-A AC Power Flow Model
+We consider a power network in which one bus is selected as the reference with voltage angle fixed at zero.
+Let $V_{i}$ and $\theta_{i}$ denote the voltage magnitude and phase angle at bus $i$, respectively.
+The active and reactive power injections are represented by $P_{i}$ and $Q_{i}$.
+Subscripts $(\cdot)_{i}$ refer to quantities associated with bus $i$, while $(\cdot)_{ik}$ indicates quantities related to the branch connecting buses $i$ and $k$.
+The steady-state AC power flow equations at bus $i$ are given by
+$P_{i}$ $=V_{i}^{2}G_{ii}+\sum_{k\in\mathcal{B}_{i}}V_{i}V_{k}\left(G_{ik}\cos\theta_{ik}+B_{ik}\sin\theta_{ik}\right),$ (1a)
+$Q_{i}$ $=-V_{i}^{2}B_{ii}+\sum_{k\in\mathcal{B}_{i}}V_{i}V_{k}\left(G_{ik}\sin\theta_{ik}-B_{ik}\cos\theta_{ik}\right),$ (1b)
+where $\theta_{ik}:=\theta_{i}-\theta_{k}$, and $G$ and $B$ denote the real and imaginary parts of the network admittance matrix.
+The set $\mathcal{B}_{i}$ contains bus $i$ and all buses directly connected to it.
+The nonlinear trigonometric and multiplicative terms in (1) introduce nonconvexities that significantly complicate optimization problems embedding AC network constraints.
+This motivates the construction of tractable linear representations of selected quantities derived from these equations.
 II-B Conservative Linear Approximations
-The nonlinearity of the power flow equations in (1) contributes to the complexity encountered in solving optimization problems.
-To address this challenge, we previously introduced a sample-based conservative linear approximation (CLA) approach aimed at either over- or under-estimating specified quantities of interest, such as the magnitudes of voltages and current flows (as illustrated in Fig.
-1)[12].
-Moreover, CLAs facilitate parallel computation by enabling concurrent computation of the CLA for each quantity of interest.
-The construction of a CLA entails sampling power injections across an operational range of interest, such as a range of loads and power generated by Distributed Energy Resources (DERs), followed by computing power flow solutions for each sample and solving a constrained-regression problem.
-For instance, samples for load demands are acquired utilizing a predefined probability distribution $\mathbb{P}_{\mathcal{S}}$ over a specified operational range $\mathcal{S}$.
-This range could be defined as $\mathcal{S}=\{P_{L_{d}}^{\text{min}}\leq P_{L_{d}}\leq P_{L_{d}}^{\text{max}},\enspace Q_{L_{d}}^{\text{min}}\leq Q_{L_{d}}\leq Q_{L_{d}}^{\text{max}}\mbox{for all }L_{d}\in\mathcal{N_{D}}\}$, where $(\,\cdot\,)_{L_{d}}$ denotes the load demand, $\mathbb{P}_{\mathcal{S}}$ represents the uniform distribution, $\mathcal{N_{D}}$ is the set of all buses with loads, and the superscripts max (min) indicate upper (lower) limits.
-Figure 1: An illustration showcasing a comparison between a conventional linear approximation (on the left) and CLAs (on the right).
-In this visual representation, the solid line signifies the nonlinear function under consideration.
-In the figure on the left, the dotted line represents a traditional first-order Taylor approximation centered at point $\times$, while in the figure on the right, the dotted line above (below) corresponds to an over- (under-)estimating approximation.
-The utilization of CLAs allows for the customization of the approximation to fit a specific operating range and the targeted system.
-Additionally, the sample-based approach enables integrating the behavior of complicated devices like tap-changing transformers and smart inverters into the approximation, as discussed in our prior work [6].
-In the realm of optimization, CLAs offer a crucial advantage: they enable the satisfaction of nonlinear constraints while enforcing only linear inequalities, assuming the CLAs maintain conservativeness.
-Consequently, CLAs streamline optimization problems, rendering them suitable for commercial optimization solvers.
-Consider a quantity of interest denoted as $\gamma$, which could represent variables such as the voltage magnitude at a specific bus or the magnitude of current flow along a particular line.
-In this context, bold quantities signify matrices and vectors.
-Let superscript $T$ denote the transpose.
-An overestimating CLA can be expressed as follows:
-$a_{0}+\bm{a}_{1}^{T}\begin{bmatrix}\bm{P}\\ \bm{Q}\end{bmatrix}$ (2)
-where $a_{0}$ is a scalar and $\bm{a}_{1}$ is a vector, both serving as decision variables in the regression problem later described in (4).
-This CLA is designed to ensure the following relationship for power injections $\bm{P}$ and $\bm{Q}$ within a specified range:
+To improve tractability while preserving feasibility with respect to the AC model, we previously developed a sample-based CLA framework [BUASON2022].
+The goal of CLA is to construct affine functions of power injections that approximate nonlinear quantities of interest—such as voltage magnitudes or current flows—while ensuring one-sided error behavior over a prescribed operating region.
+The CLA procedure begins by defining an operating set $\mathcal{S}$ for power injections. For example, load demands may be sampled from a probability distribution $\mathbb{P}_{\mathcal{S}}$ over
+$\mathcal{S}=\{P_{L_{d}}^{\min}\leq P_{L_{d}}\leq P_{L_{d}}^{\max},\;Q_{L_{d}}^{\min}\leq Q_{L_{d}}\leq Q_{L_{d}}^{\max},\;\forall L_{d}\in\mathcal{N}_{D}\},$
+where $\mathcal{N}_{D}$ denotes buses with loads and the superscripts indicate lower and upper bounds.
+A uniform distribution is commonly adopted.
+For each sampled injection vector, the AC power flow equations are solved to obtain the corresponding nonlinear quantities of interest.
+These samples are then used to compute an affine approximation through constrained regression.
+An illustrative comparison between a conventional first-order Taylor expansion and a CLA is shown in Fig.
+1.
+Unlike a local linearization centered at a single operating point, a CLA is constructed to systematically overestimate or underestimate the nonlinear function throughout the sampled region.
+Figure 1: Comparison of a conventional linearization (left) and CLAs (right).
+The solid curve represents the nonlinear function.
+The dotted line on the left corresponds to a first-order Taylor approximation at the marked point.
+The dotted lines on the right represent over- and under-estimating conservative approximations.
+In addition to improved regional fidelity, CLAs provide two practical advantages.
+First, each quantity of interest can be approximated independently, enabling parallel computation of coefficients.
+Second, if conservativeness is maintained, nonlinear inequality constraints can be replaced with linear inequalities while preserving feasibility relative to the AC model.
+This enables compatibility with commercial linear and mixed-integer solvers.
+Let $\gamma$ denote a scalar quantity derived from the AC power flow solution, such as a voltage magnitude or branch current magnitude.
+Using vector notation for power injections, an overestimating CLA takes the affine form
+$a_{0}+\bm{a}_{1}^{T}\begin{bmatrix}\bm{P}\\ \bm{Q}\end{bmatrix}$, (2)
+where $a_{0}$ and $\bm{a}_{1}$ are regression coefficients. Conservativeness requires that, for injections within the operating range,
 $\gamma\leq a_{0}+\bm{a}_{1}^{T}\begin{bmatrix}\bm{P}\\ \bm{Q}\end{bmatrix}$. (3)
-Assuming that (3) is indeed satisfied, we can ensure that the constraint $\gamma\leq\gamma^{\text{max}}$ is also satisfied by instead enforcing a linear constraint $a_{0}+\bm{a}_{1}^{T}\begin{bmatrix}\bm{P}\\ \bm{Q}\end{bmatrix}\leq\gamma^{\text{max}}$.
-This approach allows us to meet the upper bound requirement $\gamma^{\text{max}}$ without introducing the implicit system of nonlinear AC power flow equations in (1).
-Importantly, by employing the CLA, we are able to satisfy the nonlinear equations while maintaining a linear formulation, thus enhancing computational tractability without sacrificing feasibility in the resulting solution.
-To compute a CLA, we solve for the coefficients of the affine function of power injections in (2) in the following regression problem:
-$\min_{a_{0},\ \bm{a}_{1}}\quad\frac{1}{M}\sum_{m=1}^{M}\mathcal{L}\left(\gamma_{m}-\left(a_{0}+\bm{a}_{1}^{T}\begin{bmatrix}\bm{P}_{m}\\ \bm{Q}_{m}\end{bmatrix}\right)\right)$ (4a)
+If (3) holds, then an upper bound constraint $\gamma\leq\gamma^{\max}$ can be enforced through the linear condition
+$a_{0}+\bm{a}_{1}^{T}\begin{bmatrix}\bm{P}\\ \bm{Q}\end{bmatrix}\leq\gamma^{\max},$
+without explicitly embedding the nonlinear AC equations. Thus, feasibility with respect to the nonlinear model is preserved while maintaining linear structure.
+The coefficients in (2) are obtained by solving the constrained regression problem
 […]
-The subscript $(\,\cdot\,)_{m}$ denotes the $m^{\text{th}}$ sample and $M$ is the number of samples.
-The function $\mathcal{L}(\,\cdot\,)$ represents a loss function, such as the absolute value for $\ell_{1}$ loss or the square for squared-$\ell_{2}$ loss.
-In this paper, our focus is on quantities of interest denoted by $\gamma$, which correspond to the magnitudes of voltages ($V$) and current flows ($I$).
-The construction of underestimating CLAs follows a similar process as described in (4), with the key distinction being the reversal of the inequality direction in (4b).
-The conservativeness of the CLA computed in (4) comes at the cost of reduced accuracy relative to the approximation corresponding to the unconstrained regression problem resulting from dropping (4a) from (4).
-To manage this tradeoff, the next section presents the main contribution of this paper, namely, a linear approximation technique that achieves a balance between conservativeness and accuracy.
-This approach involves biasing the linearization towards conservativeness, guided by a designated loss function.
+s.t. $\gamma_{m}-\left(a_{0}+\bm{a}_{1}^{T}\begin{bmatrix}\bm{P}_{m}\\ \bm{Q}_{m}\end{bmatrix}\right)\leq 0,\quad m=1,\dots,M.$ (4b)
+where $m$ indexes samples, $M$ denotes the number of samples, and $\mathcal{L}(\cdot)$ is a loss function such as the absolute value ($\ell_{1}$) or squared error ($\ell_{2}$).
+Underestimating CLAs are obtained by reversing the inequality in (4b).
+While the constraint in (4b) guarantees directional conservativeness, it restricts the feasible regression set and may increase approximation error compared to the unconstrained formulation obtained by removing this condition.
+Consequently, strict conservativeness can reduce overall accuracy, particularly in regions where the nonlinear relationship is difficult to capture with a single affine model.
+The next section introduces the main contribution of this paper, which extends the CLA framework by incorporating directional bias directly into the regression objective.
+This formulation enables a more flexible trade-off between feasibility protection and approximation accuracy.
 III Conservative Bias Linear Approximations
-The CLA approach presented in Section II-B is consistently conservative within the set of drawn samples.
-However, in certain scenarios, the conservativeness property may lead to significant errors due to specific samples.
-In this paper, we present a sample-based conservative bias linear approximation (CBLA) approach that is adaptive, meaning it can be tailored to a specific system and operating range.
-The CBLA approach is designed to be optimal, minimizing a specific error metric while maintaining a tendency to be conservative to enhance accuracy.
-This means CBLA focuses on minimizing errors between the approximating function and the samples, allowing some samples to violate conservativeness at a specified cost.
+The CLA framework described in Section II-B guarantees strict one-sided conservativeness across all sampled operating points.
+While this property is desirable for feasibility protection, enforcing it uniformly may lead to unnecessarily large approximation errors when certain samples are difficult to represent accurately with a single affine model.
+To address this limitation, we introduce a sample-based conservative bias linear approximation (CBLA) methodology.
+Similar to CLA, the proposed approach is adaptive in that it is constructed using samples drawn from a specified operating region for a particular system.
+However, instead of imposing conservativeness as a hard constraint, CBLA incorporates a bias toward conservativeness directly into the optimization objective.
+This formulation enables the approximation to remain predominantly conservative while allowing limited violations at a controlled cost.
+As a result, CBLA seeks to achieve a principled balance between directional bias and approximation accuracy.
 III-A Formulation
-Let $\epsilon$ denote the mismatch between the approximated quantity and the actual quantity. The optimization problem to compute a CBLA is formulated as follows:
-$\min_{f(\epsilon_{m}(a_{0},\bm{a}_{1}))}\quad\frac{1}{M}\sum_{m=1}^{M}f(\epsilon_{m})$ (5)
+Let $\epsilon_{m}$ denote the approximation error (mismatch) at sample $m$, defined as the difference between the true quantity of interest and its affine approximation.
+The CBLA coefficients are computed by solving the optimization problem
+$\min_{a_{0},\,\bm{a}_{1}}\quad\frac{1}{M}\sum_{m=1}^{M}f(\epsilon_{m}),$ (5)
 where
 $[\forall m=1,\ldots,M]$
-$\epsilon_{m}=\gamma_{m}-\left(a_{0}+\bm{a}_{1}^{T}\begin{bmatrix}\bm{P}_{m}\\ \bm{Q}_{m}\end{bmatrix}\right),$ (6)
-and
-$f(\epsilon_{m})$ =$=\begin{dcases}g(\epsilon_{m}),&\text{if}\;\epsilon_{m}\leq 0\\ h(\epsilon_{m}),&\text{otherwise.}\\ \end{dcases}$ (7)
-The optimization problem in (5) seeks to minimize the aggregated value of the error function $f(\,\cdot\,)$ defined in (7) over all samples by computing the coefficients $a_{0}$ and $\bm{a_{1}}$ in (6).
-This error function is contingent upon the error mismatch, denoted as $\epsilon$ in (6), between the estimated quantity and the actual quantity ($\gamma$).
-The error function is computed based on the error’s sign for each sample.
-For overestimation, $h(\epsilon_{m})$ is set high to impose a substantial cost for violation, while $g(\epsilon_{m})$ is kept low for non-violations.
-Conversely, for underestimation, $h(\epsilon_{m})$ is relatively low and $g(\epsilon_{m})$ is high.
-For an alternative formulation of CBLA that incorporates a penalty term to influence the optimization behavior, see the appendix.
-III-B Error function
-Choosing a suitable error function in the CBLA approach is an important consideration.
-The choice of error function depends on various factors, such as the specific system requirements, a quantity of interest, and the trade-off between accuracy and conservativeness.
-To better understand how the error function works, we compare the error function used in the CLA approach with that of the CBLA approach.
-The CLA approach imposes conservativeness across the set of sampled data in the constraints.
-We can rewrite the regression problem described in (4), which utilizes the $\ell_{1}$ loss function, as an optimization problem formulated in (5)–(7).
-In this formulation, the error function is defined as follows:
-$f(\epsilon_{m})$ =$=\begin{dcases}\epsilon_{m},&\text{if}\;\epsilon_{m}\leq 0\\ \infty,&\text{otherwise.}\\ \end{dcases}$ (8)
-Figure 2: An error function of the CLA where $g(\epsilon)=\epsilon$ and $h(\epsilon)=\infty$.
-The error function in (8) (see Fig. 2) assigns an infinite cost to any violation of the overestimating requirement. This implies that for all samples drawn, no violation is permitted.
-Figure 3: An example of an error function where $g(\epsilon)=\epsilon$ and $h(\epsilon)=e^{2\epsilon}-1$.
-In contrast, our CBLA approach offers the flexibility to configure an error function that accommodates violations for specific samples, all while considering a predefined cost associated with these violations.
-This ability to tailor the error function empowers us to strike a balance between accuracy and the acceptable level of conservativeness for the specific system and operating range of interest.
-In Fig.
-3, an example of an error function $f(\epsilon)$ for an overestimating CBLA is depicted.
-The function $f(\epsilon)$ exhibits a high value for $\epsilon>0$, $f(\epsilon)$ maintains a relatively low value for $\epsilon<0$, and $f(0)$ is zero, indicating an exact approximation.
-In this setup, assigning a higher penalty for violations of conservativeness leads to more conservative approximations at the expense of some accuracy.
-Additionally, when the derivative of the function $h(\epsilon)$ increases, as shown in Fig.
-3, the function tends to allow only small positive values of $\epsilon$, because larger positive values result in exponentially higher penalties.
-Conversely, reducing the penalty cost tends to produce more accurate approximations but with less conservativeness.
-When the error function is defined as a piecewise linear function (i.e., when both $g(\epsilon)$ and $h(\epsilon)$ are linear functions), the problem formulation to compute CBLAs in (5)–(7) can be framed as a linear program as follows:
+[…]
+The function $f(\epsilon_{m})$ determines how errors are penalized and is defined piecewise as
+$f(\epsilon_{m})=\begin{dcases}g(\epsilon_{m}),&\text{if }\epsilon_{m}\leq 0,\\ h(\epsilon_{m}),&\text{otherwise.}\end{dcases}$ (7)
+The optimization problem in (5) therefore minimizes the average penalized error across all samples by selecting coefficients $a_{0}$ and $\bm{a}_{1}$.
+The key distinction from CLA lies in the structure of the error function.
+Rather than forbidding violations of conservativeness, CBLA assigns asymmetric penalties based on the sign of $\epsilon_{m}$.
+For an overestimating approximation, violations correspond to $\epsilon_{m}>0$.
+In this case, $h(\epsilon_{m})$ is chosen to impose a relatively large penalty, while $g(\epsilon_{m})$ remains comparatively small for non-violating samples.
+For underestimating approximations, the roles of $g(\cdot)$ and $h(\cdot)$ are reversed.
+This asymmetric design induces a tendency toward conservativeness without strictly enforcing it.
+III-B Error Function Design
+The selection of the error function $f(\cdot)$ plays a central role in determining the trade-off between conservativeness and approximation accuracy.
+Its design may depend on system requirements, the specific quantity being approximated, and the acceptable level of directional violation.
+To illustrate the relationship between CLA and CBLA, consider the CLA regression problem in (4) with an $\ell_{1}$ loss. That formulation can be interpreted within the CBLA framework by defining
+$f(\epsilon_{m})=\begin{dcases}\epsilon_{m},&\text{if }\epsilon_{m}\leq 0,\\ \infty,&\text{otherwise.}\end{dcases}$ (8)
+Figure 2: Error function corresponding to CLA, where $g(\epsilon)=\epsilon$ and $h(\epsilon)=\infty$.
+As shown in Fig.
+2, assigning infinite cost to $\epsilon_{m}>0$ strictly prohibits violations, thereby recovering the CLA formulation.
+In contrast, CBLA replaces the infinite penalty with a finite but potentially large function, permitting limited violations when doing so improves overall accuracy.
+Figure 3: Example CBLA error function for an overestimating approximation, where $g(\epsilon)=\epsilon$ and $h(\epsilon)=e^{2\epsilon}-1$.
+An example is shown in Fig.
+3, where $g(\epsilon)=\epsilon$ and $h(\epsilon)=e^{2\epsilon}-1$.
+In this case, positive errors incur exponentially increasing penalties, discouraging large violations while allowing small deviations when necessary.
+Increasing the growth rate of $h(\epsilon)$ promotes stronger conservativeness, whereas reducing its slope yields a more accurate but less conservative approximation.
+When both $g(\epsilon)$ and $h(\epsilon)$ are linear functions, i.e., the error function is piecewise linear, the CBLA regression problem can be reformulated as a linear program:
 $\min_{a_{0},\,\bm{a}_{1}}\quad$ $\frac{1}{M}\sum_{m=1}^{M}z_{m}$ (9a)
+s.t. $[\forall m=1,\ldots,M]$
 […]
-[…]
-[…]
-[…]
-where $z$ is a slack variable, and $k_{1}$ and $k_{2}$ are the coefficients of the linear error functions, i.e., $g(\epsilon)=k_{1}\epsilon$ and $h(\epsilon)=k_{2}\epsilon$.
-Nonlinear error functions transform the regression problem into a mixed-integer nonlinear program.
-This can be efficiently implemented using a user-defined function in Julia and solved with packages such as Optim, which offers a robust framework for constrained optimization problems [20].
+$z_{m}\geq k_{1}\epsilon_{m},$ (9c)
+$z_{m}\geq k_{2}\epsilon_{m},$ (9d)
+where $z_{m}$ is an auxiliary variable and $k_{1}$, $k_{2}$ define the slopes of $g(\epsilon)=k_{1}\epsilon$ and $h(\epsilon)=k_{2}\epsilon$, respectively.
+If nonlinear penalty functions are selected, the resulting optimization problem becomes a nonlinear program (or potentially a mixed-integer nonlinear program, depending on the formulation).
+Such problems can be implemented using user-defined functions in Julia and solved using packages such as Optim, which provides a flexible framework for constrained optimization [optimjl].
 IV Unit Commitment
@@
 The primary goal is to minimize the total cost of operation while ensuring that the demand for electricity is met and the technical and operational constraints of the generation units and transmission network are satisfied.
-The variability and uncertainty associated with loads, particularly when load shedding may be necessary, and generation introduce additional complexities that the unit commitment (UC) problem must address to ensure grid reliability and operational efficiency.
+The variability and uncertainty associated with loads, particularly when load shedding may be necessary, and generation introduce additional complexities that the UC problem must address to ensure grid reliability and operational efficiency.
 As a mixed-integer optimization problem, the UC formulation involves binary variables to represent the on/off status of generators, along with continuous variables to model power output.
 IV-A Computational Procedure and Parallelization Potential
-Figure 4: A flowchart illustrating the computational process for solving the unit commitment (UC) problem using CBLA.
+Figure 4: A flowchart illustrating the computational process for solving the UC problem using CBLA.
 Steps enclosed in red boxes correspond to CBLA, while steps enclosed in blue boxes relate to the UC problem.
 Steps marked with * indicate parallelizable processes.
-Solving the unit commitment problem using CBLA involves multiple computational steps, each contributing to accurately determining generator commitments and dispatch while leveraging efficient approximations of power flow equations.
-Given the complexity of unit commitment and the nonlinearity of AC power flow, CBLA serves as an effective method to construct computationally efficient linear approximations.
+Solving the UC problem using CBLA involves multiple computational steps, each contributing to accurately determining generator commitments and dispatch while leveraging efficient approximations of power flow equations.
+Given the complexity of UC and the nonlinearity of AC power flow, CBLA serves as an effective method to construct computationally efficient linear approximations.
 The overall process consists of three key steps:
@@
   2. 2.
-    Unit commitment optimization: Using the CBLA approximations obtained in Step 1, solve the UC problem to determine generator statuses for each generator at each hour while ensuring operational constraints are satisfied.
+    UC optimization: Using the CBLA approximations obtained in Step 1, solve the UC problem to determine generator statuses for each generator at each hour while ensuring operational constraints are satisfied.
   3. 3.
-    AC optimal power flow (AC OPF) evaluation: Given the generator statuses from Step 2, solve the AC OPF that allows load shedding to determine the generation dispatch at each hour, ensuring feasibility under AC power flow constraints [21].
+    AC optimal power flow (AC OPF) evaluation: Given the generator statuses from Step 2, solve the AC OPF that allows load shedding to determine the generation dispatch at each hour, ensuring feasibility under AC power flow constraints [Happ1977].
 This computational framework allows parallel execution, particularly in Steps 1 and 3.
@@
 Fig.
-4 illustrates the full set of computational steps required to solve the unit commitment problem using the CBLA approach.
+4 illustrates the full set of computational steps required to solve the UC problem using the CBLA approach.
 IV-B Optimization Problem Formulation
 Let $u_{g,t}$ be a generator on/off status and $s_{g,t}\;(r_{g,t})$ be a generator startup (shutdown) status for generator $g$ at time $t$.
-The unit commitment problem with the DC power flow [22] is formulated as follows:
-$\min$ $\sum_{t\in\mathcal{T}}\left(\sum_{g\in\mathcal{G}}c_{g}(P_{g,t})+C_{g}^{\text{startup}}s_{g,t}+C_{g}^{\text{shutdown}}r_{g,t}\right)$ (10a)
+The UC problem with the DC power flow [vandenbergh2014] is formulated as follows:
+$\min\$$\sum_{t\in\mathcal{T}}\left(\sum_{g\in\mathcal{G}}c_{g}(P_{g,t})+C_{g}^{\text{startup}}s_{g,t}+C_{g}^{\text{shutdown}}r_{g,t}\right)$ (10a)
 s.t. $\sum_{g\in\mathcal{G}}P_{g,t}=D_{t},\quad\forall t\in\mathcal{T},$ (10b)
@@
 IV-C Unit Commitment using Conservative Bias Linear Approximations
-The unit commitment problem formulated in the previous subsection enforces network feasibility using a DC power flow approximation.
+The UC problem formulated in the previous subsection enforces network feasibility using a DC power flow approximation.
 While computationally efficient, the DC power flow model does not explicitly incorporate voltage and current constraints.
@@
 4).
-In each iteration, we solve the UC problem with the current CBLA, then check if the generation statuses and dispatch levels match those from a first-order Taylor-based method or previous CBLA iterations.
-We also verify whether power injections stay within the sampled region used to build the CBLA.
-If violations occur, we identify the affected time periods and buses or lines, draw new samples around the updated solution, recompute the CBLA, and resolve the UC problem.
-This process prevents accuracy degradation from extrapolation and cycling between inconsistent UC solutions.
+In each iteration, the UC problem is solved using the current CBLA, and the resulting generation statuses and dispatch levels are checked against those from a first-order Taylor-based method or previous CBLA iterations.
+We also verify that power injections remain within the sampled region used to construct the CBLA.
+If violations are detected, new samples are drawn around the updated solution, the CBLA is recomputed, and the UC problem is resolved.
+This process mitigates extrapolation errors and prevents cycling between inconsistent UC solutions.
+In addition, the search can be restricted to a limited set of candidate UC solutions generated by the solver.
+For instance, a small number of feasible candidates (e.g., up to 20) may be evaluated and compared based on feasibility and operational cost, with each candidate validated through an AC OPF evaluation.
+This approach provides a practical trade-off between solution quality and computational effort by selecting the best available candidate within the allowed computation time.
 This CBLA-based formulation further allows two practical enhancements.
-First, it offers flexibility in controlling conservativeness by adjusting a bias term when constructing the constraints, enabling the user to trade off between feasibility margin and conservatism.
+First, it offers flexibility in controlling conservativeness by adjusting a bias term when constructing the constraints, enabling the user to trade off between feasibility margin and conservativeness.
 Second, it supports the inclusion of penalty terms that allow for controlled constraint violations, enhancing feasibility in cases where a strictly conservative formulation may be overly restrictive.
 IV-D Combinatorial Growth of Valid Generator Statuses
-In the unit commitment formulation, each generator’s on/off status over time must satisfy minimum up and down time constraints (10e)–(10f), substantially reducing the number of admissible on/off sequences.
+In the UC formulation, each generator’s on/off status over time must satisfy minimum up and down time constraints (10e)–(10f), substantially reducing the number of admissible on/off sequences.
 Nevertheless, the feasible status space still grows exponentially with the number of time periods, making brute-force enumeration impractical.
@@
 The recurrence relation is:
-$f(\lvert\mathcal{T}\rvert,s,\ell)=\begin{cases}\sum\limits_{k=d}^{\lfloor\nicefrac{{\lvert\mathcal{T}\rvert}}{{2}}\rfloor}f(\lvert\mathcal{T}\rvert-k,1-s,k),\;\text{if }\ell=d,\lvert\mathcal{T}\rvert\geq d\\[4.30554pt] f(\lvert\mathcal{T}\rvert-1,s,\ell-1),\;\text{if }\ell>d\\[4.30554pt] 1,\;\text{if }\lvert\mathcal{T}\rvert=\ell\leq d\text{, one of the exceptions holds}\\[4.30554pt] 0,\;\text{otherwise}\end{cases}$
+[…]
 where the exceptions include:
@@
 For example, with $\lvert\mathcal{T}\rvert=5$ and $d=2$, applying the recurrence yields 13 valid sequences for a single generator.
-Therefore, for 6 generators, the total number of configurations is $(F(5))^{6}=13^{6}=4,\!826,\!809$, demonstrating how the feasible status space grows rapidly even for modest systems.
+Therefore, for six generators, the total number of configurations is $(F(5))^{6}=13^{6}=4,\!826,\!809$, demonstrating how the feasible status space grows rapidly even for modest systems.
 This growth illustrates why brute-force enumeration of all valid generator commitment combinations quickly becomes intractable.
-To manage this complexity, we avoid exhaustive search by solving the unit commitment problem iteratively: each iteration only considers a manageable subset of decisions, and additional constraints are introduced in subsequent iterations to guide the solution away from unstable commitment patterns, such as cycling or oscillating generator decisions, i.e., repeatedly looping through statuses.
+To manage this complexity, we avoid exhaustive search by solving the UC problem iteratively: each iteration only considers a manageable subset of decisions, and additional constraints are introduced in subsequent iterations to guide the solution away from unstable commitment patterns, such as cycling or oscillating generator decisions, i.e., repeatedly looping through statuses.
 The algorithm is explained in more detail in Section IV-A.
@@
 The AC Optimal Power Flow (AC OPF) problem determines the optimal generation dispatch by solving the full nonlinear AC power flow equations while satisfying network constraints.
-Using generator commitment statuses from the unit commitment problem in (10) (solved with formulations such as DC power flow, first-order Taylor approximations, or CBLAs), we compute the dispatch for each online generator at each hour.
+Using generator commitment statuses from the UC problem in (10) (solved with formulations such as DC power flow, first-order Taylor approximations, or CBLAs), we compute the dispatch for each online generator at each hour.
 The multi-period formulation minimizes the total cost across all hours while ensuring power balance and network feasibility at each time.
@@
 The AC OPF with load shedding is formulated as follows:
-$\min$ $\sum_{t\in\mathcal{T}}\left(\sum_{g\in\mathcal{G}}c_{g}(P_{g,t})+\sum_{i\in\mathcal{B}}C^{\text{shed}}(L_{i,t})\right)$ (14a)
+$\min\$$\sum_{t\in\mathcal{T}}\left(\sum_{g\in\mathcal{G}}c_{g}(P_{g,t})+\sum_{i\in\mathcal{B}}C^{\text{shed}}(L_{i,t})\right)$ (14a)
 s.t. $[\forall i\in\mathcal{B},\forall t\in\mathcal{T},\forall g\in\mathcal{G},\forall(i,k)\in\mathcal{L}]$
@@
 The power balance constraints in (14b)–(14c) enforce the nonlinear AC power flow equations for active and reactive power at each bus and time period.
-The generator limits in (14d)–(14e) ensure that active and reactive power generation respects the unit commitment status $u_{g,t}$.
+The generator limits in (14d)–(14e) ensure that active and reactive power generation respects the UC status $u_{g,t}$.
 The voltage limits in (14f) ensure that bus voltages remain within safe operating ranges.
@@
 Finally, the load shedding constraints in (14h) ensure that the curtailed demand remains non-negative and does not exceed the original load at any bus and time period.
-This formulation integrates the unit commitment results into AC OPF, optimizing the dispatch of committed generators while allowing for load shedding with constant power factors.
+This formulation integrates the UC results into AC OPF, optimizing the dispatch of committed generators while allowing for load shedding with constant power factors.
 V Numerical results
 In this section, we conduct numerical experiments on several test cases to examine the behavior of CBLA, highlight the benefits of error function design, and demonstrate the effectiveness of CBLA in a UC problem.
-The test cases used in the simulations are the IEEE 24-bus system and the IEEE 30-bus system, all of which are accessible in Matpower [23].
+The test cases used in the simulations are the IEEE 24-bus system and the IEEE 30-bus system, both of which are accessible in Matpower [zimmerman_matpower_2011].
 For approximations of voltage and current flow, we draw $500$ samples by varying the power injections within a range of $70\%$ to $130\%$ of their nominal values.
@@
 We use the $\ell_{1}$ norm as the loss function $\mathcal{L}(\,\cdot\,)$.
-The numerical simulation was conducted in MATLAB using Yalmip [24] and in Julia using the Optim package.
+The numerical simulation was conducted in MATLAB using Yalmip [Lofberg2004] and in Julia using the Optim package.
+The number of samples (500–1000 in the reported experiments) is selected to provide adequate coverage of the multidimensional injection space while maintaining computational tractability.
+Empirically, we observe that beyond this range (e.g., more than 500–1000 samples for the tested systems), the marginal improvement in approximation accuracy and solution robustness becomes negligible relative to the additional computational cost.
 V-A Conservative Bias Linear Approximations
@@
 V-B Application: Unit Commitment
-The comparison is conducted on the IEEE 30-bus test system, which has 6 generators.
+The comparison is conducted on the IEEE 30-bus test system, which has six generators.
 Voltage magnitude limits are set to 0.94 p.u.
@@
 TABLE II: Hourly Comparison Across Different Approximations
-H1 H2 H3 H4 H5
+Hour
+1 2 3 4 5
 Load Demand (MW) 152.2 192.13 234.29 287.33 219.27
@@
 Gen 3 – Gen 6 0 0 0 0 0
-Total Cost (\$) 1818 6328 6090 20417 18869
+Total Cost (\$) 1817.6 6328 6090 20417 18869
 Gen-Only Cost (\$) 1817.6 2135.2 3076.1 3431.5 2167.3
@@
 This trade-off underscores the need to balance conservativeness and approximation accuracy.
-CBLA’s tunable parameter $\alpha$ allows us to adjust this balance to achieve the desired feasibility and flexibility.
+CBLA’s tunable parameter $\alpha$ allows for adjusting this balance to achieve the desired level of feasibility robustness and operational flexibility.
 By tailoring the approximation to system conditions, CBLA helps reduce constraint violations and yields UC solutions that generally remain feasible under the nonlinear AC power flow model.
+A practical approach to selecting $\alpha$ is to evaluate constraint violations on an independent set of AC-feasible operating points and adjust the parameter accordingly.
+If $\alpha$ is chosen too small, the approximation may become insufficiently conservative, potentially leading to AC infeasibility in the UC solution.
+Conversely, excessively large values of $\alpha$ result in a more conservative formulation, which preserves feasibility but may increase operational cost.
+In either case, the impact primarily affects the optimality–conservativeness balance rather than the structural validity of the method.
 Moreover, the benefits of CBLA become more pronounced in larger systems or under tight feasibility margins.
 Unlike the first-order Taylor approximation, which is inherently local, CBLA is constructed to remain conservative across a sampled region of interest.
-This range-aware property improves cost-efficiency in decision-making.
-Compared to traditional methods, CBLA provides a more adaptive solution framework that maintains the tractability of linear programming while reducing the gap between the models of linear and nonlinear power systems.
+This range-aware property improves cost efficiency.
+Compared to traditional methods, CBLA provides a more adaptive solution framework that maintains the tractability of linear programming while reducing the gap between linear and nonlinear power system models.
 VI Conclusion and Future Work
@@
 Additionally, the ability to choose different error functions allows CBLA to be tailored to specific systems and operational conditions.
-When applied to the unit commitment problem, the CBLA-based formulation consistently outperformed formulations based on DC power flow and first-order Taylor approximations, achieving lower overall operating costs in the subsequent evaluation of the AC optimal power flow.
-In our future work, we aim to extend our current approach by developing additional conservative bias approximations through the use of piecewise linearizations.
+When applied to the UC problem, the CBLA-based formulation consistently outperformed formulations based on DC power flow and first-order Taylor approximations, achieving lower overall operating costs in the subsequent evaluation of the AC optimal power flow.
+For benchmarking purposes, we compared CBLA against the DC power flow model and the first-order Taylor approximation, as these remain among the most widely adopted linearization approaches in both industry practice and academic research.
+In particular, DC power flow is the standard approximation for large-scale UC formulations, while Taylor-based linearizations are commonly used in optimization-based power system studies.
+A systematic comparison with other advanced linearized AC models (e.g., LPAC [Coffrin2014] and other formulations [fnt]) is a topic for future work.
+In addition, our future work aims to extend our current approach by developing additional CBLA through the use of piecewise linearizations.
 Moreover, we plan to apply our proposed approach to a broader range of power system planning and resilience tasks.
@@
 Appendix: First-Order Taylor Approximation of Voltage and Current Constraints
-Suppose the power injections $\bm{P}$ and $\bm{Q}$ deviate from their nominal values $\bm{P}^{\text{nom}}$ and $\bm{Q}^{\text{nom}}$.
-Let $\mathcal{B}^{\text{PQ}}$ denote the set of PQ buses, $\widetilde{V}_{i,t}$ denote the first-order Taylor approximation of the voltage magnitude at bus $i$ and time $t$, and let the abbreviation “nom” indicate that a quantity is evaluated at the nominal operating point, which is computed separately for each hour based on the corresponding loads.
-For each $i\in\mathcal{B}^{\text{PQ}}$, the first-order Taylor approximation of the voltage magnitude is:
-$\widetilde{V}_{i,t}=V_{i}^{\text{nom}}+\frac{\partial V_{i}}{\partial\bm{P}}\bigg{|}_{\text{nom}}\cdot(\bm{P}_{t}-\bm{P}^{\text{nom}})+\frac{\partial V_{i}}{\partial\bm{Q}}\bigg{|}_{\text{nom}}\cdot(\bm{Q}_{t}-\bm{Q}^{\text{nom}}),$ (15)
-where the partial derivatives $\frac{\partial V_{i}}{\partial\bm{P}}$ and $\frac{\partial V_{i}}{\partial\bm{Q}}$ are evaluated at the nominal operating point using the inverse of the power flow Jacobian.
-For non-PQ buses, the voltage magnitude is fixed at its nominal value:
+Let $\bm{P}$ and $\bm{Q}$ denote the vectors of active and reactive power injections, respectively, and let $\bm{P}^{\text{nom}}$ and $\bm{Q}^{\text{nom}}$ represent their corresponding nominal values.
+The nominal operating point is determined independently for each time period according to the associated load conditions.
+Let $\mathcal{B}^{\text{PQ}}$ denote the set of PQ buses, and let $\widetilde{V}_{i,t}$ represent the linearized estimate of the voltage magnitude at bus $i$ during time period $t$.
+The superscript “nom” indicates that a quantity is evaluated at the nominal operating point.
+For buses $i\in\mathcal{B}^{\text{PQ}}$, the voltage magnitude is approximated by a first-order expansion around the nominal operating point:
+$\widetilde{V}_{i,t}=V_{i}^{\text{nom}}+\left.\frac{\partial V_{i}}{\partial\bm{P}}\right|_{\text{nom}}\cdot(\bm{P}_{t}-\bm{P}^{\text{nom}})+\left.\frac{\partial V_{i}}{\partial\bm{Q}}\right|_{\text{nom}}\cdot(\bm{Q}_{t}-\bm{Q}^{\text{nom}}),$ (15)
+where the sensitivity matrices $\frac{\partial V_{i}}{\partial\bm{P}}$ and $\frac{\partial V_{i}}{\partial\bm{Q}}$ are evaluated at the nominal operating point using the inverse of the power flow Jacobian.
+For buses outside the PQ set, the voltage magnitude is assumed to remain fixed at its nominal value:
 $\widetilde{V}_{i,t}=V_{i}^{\text{nom}},\quad\forall i\notin\mathcal{B}^{\text{PQ}}.$ (16)
-To model current flow limits, we approximate the real part of the power flow on each transmission line $(i,k)\in\mathcal{L}$ using the difference in bus voltage magnitudes from (16):
-$\Re(S_{ik,t})\approx|Y_{ik}|\cdot(\widetilde{V}_{i,t}-\widetilde{V}_{k,t}),$ (17)
-where $\Re(\;\cdot\;)$ denotes the real part of a complex quantity and $|Y_{ik}|$ is the magnitude of the admittance between buses $i$ and $k$.
-We then enforce the following absolute value inequality to approximate current magnitude limits:
-$-I_{ik}^{\text{max}}\leq|Y_{ik}|\cdot(\widetilde{V}_{i,t}-\widetilde{V}_{k,t})\leq I_{ik}^{\text{max}},\;\forall(i,k)\in\mathcal{L},\;\forall t\in\mathcal{T}.$ (18)
-To allow controlled violations of these approximations similar to (11)–(12), slack variables may be added as needed in the optimization formulation.
+To approximate current limits, the real component of the apparent power flow on each line $(i,k)\in\mathcal{L}$ is expressed as a function of the difference between the linearized voltage magnitudes obtained from (16):
+$\Re(S_{ik,t})\approx|Y_{ik}|\,(\widetilde{V}_{i,t}-\widetilde{V}_{k,t}),$ (17)
+where $\Re(\cdot)$ denotes the real part of a complex quantity and $|Y_{ik}|$ is the magnitude of the line admittance.
+Using this relationship, limits on the current magnitude can be approximated through the following linear constraints:
+$I_{ik}^{\text{max}}\leq|Y_{ik}|(\widetilde{V}_{i,t}-\widetilde{V}_{k,t})\leq I_{ik}^{\text{max}},\quad\forall(i,k)\in\mathcal{L},\;\forall t\in\mathcal{T}.$ (18)
+Finally, to provide flexibility in the optimization model, slack variables can be incorporated to allow controlled violations of the linearized constraints, similar to the treatment used in (11)–(12).